### v1

# Spatio-Temporal Graphical Counterfactuals: An Overview

Abstract
Counterfactual thinking is a critical yet challenging topic for artificial intelligence to learn knowledge from data and ultimately improve their performances for new scenarios. Many research works, including Potential Outcome Model and Structural Causal Model, have been proposed to realize it. However, their modelings, theoretical foundations and application approaches are usually different. Moreover, there is a lack of graphical approach to infer spatio-temporal counterfactuals, that considers spatial and temporal interactions between multiple units. Thus, in this work, our aim is to investigate a survey to compare and discuss different counterfactual models, theories and approaches, and further build a unified graphical causal frameworks to infer the spatio-temporal counterfactuals.
Index Terms:
Counterfactual inference, spatio-temporal graphical model, Potential Outcome Model, Structural Causal Model, causality.
I Introduction
How to enable an intelligent machine to think and answer a counterfactual question? E.g., sometimes we want to know, if a different investment strategy had been implemented, would we obtain higher returns? Or, in computer networks, what changes would occur in network load if a node’s configuration had never been changed? Or, would someone still purchase the product even if they had never been shown the advertisement? These questions generally have the counterfactual pattern, that is, “Observed $\dots$, if had done $\dots$, how would $\dots$ go?”. Thus, according to the pattern, firstly, time is necessary to consider, because the questions are queried now, but the counterfactual actions are supposed to be done in the past. Secondly, counterfactual outcomes are unobservable. The outcomes of some actions are observed if the actions have been done to the real-world system, thus these outcomes are factual. But the counterfactual outcomes are not observed, and they are imaged through supposing different actions had been done.
Figure 1: Infer counterfactuals with less and less knowledge.
To enable an intelligent machine to imitate counterfactual thinking like human, there are mainly three approaches, as shown in Fig. 1. The first is system identification, which aims to identify the system dynamics, and then solve the evolution trajectories with different initial values [1]. This process may be data-driven and adaptive, but the physical mechanisms and models are necessary to be known as knowledge in prior [2, 3, 4, 5]. E.g., in [6, 7], functional basis or network structure are necessary to be known at least one to identify the network dynamics sparsely. However, the knowledge of system mechanism is usually hard to obtain in prior.
The second approach is generative modeling, which aims to model and learn the system uncertainty, and then generate scenarios with different random inputs [8, 9]. Thus, compared with system identification, generative approach is not necessary to know the determined part, that is the knowledge of system mechanism, but on contrary, it focuses on the uncertain modeling, especially with the deep learning [10, 11, 12, 13, 14, 15]. However, the generated scenarios would not be out of the distribution of the observational data [9], while the counterfactual scenarios are usually not in the observational factuals. Thus, the mechanism knowledge is still needed to correct the bias between the distributions of factuals and counterfactuals.
Thus, the third approach is causal inference, which aims to discover causality from the observational data, and then infer factuals and counterfactuals directly with as little knowledge as possible [16, 17]. Causal counterfactual inference enables artificial intelligence to learn the knowledge from data, and ultimately improve their performances for new scenarios. To realize it, many research works have been proposed to infer counterfactuals, including the temporal cases [18, 19, 20, 21, 22, 23]. However, there are still two challenging but critical issues to address: (i) Most current inference approaches are built on the frameworks of Potential Outcome Model (POM) [17] or Structural Causal Model (SCM) [16, 24, 25, 26]. The goals of them are the same, but their foundations are different. This induces difficulty in discussing them in a unified framework. (ii) The two frameworks both cannot infer spatio-temporal counterfactuals, that considers spatial and temporal interactions between system units.
Thus, the main contributions of this work are as follows:
  1. 1.
    A survey is investigated to discuss the frameworks from POM to SCM, including their theoretical foundations and application approaches. This part is organized from Section II to Section IV.
  2. 2.
    A Forward Counterfactual Inference algorithm is designed here to recursively and autonomously infer counterfactuals through causal graphical languages. This part is presented as Algorithm 1 in Section IV-C.
  3. 3.
    An overview of the spatio-temporal graphical counterfactual framework is proposed to discuss Spatio-Temporal Bayesian Networks, its nonstationarity, and the relationship with complex networks. This part is organized in Section V.
II Potential Outcome Model
The counterfactuals are defined as the unobserved potential outcomes in POM, as shown in TABLE I. Here, $Y$ is the potential outcomes of the target variable. $T$ is treatment variable. $T=1$ represents being treated and $T=0$ represents being controlled, with a group of randomized control trials on $N$ units. $X$ represents covariates that have relationship with $Y$ and $T$, and are controlled for randomized trials. POM models the general process from causal estimands ($T$, $Y$ and $X$) to “Science” [27]. In this case, for different treatments $T$, only one potential outcome, $Y(1)$ or $Y(0)$, can be observed, and another one is unobserved, namely the counterfactual.
TABLE I: Data for POM example.
Units Covariates Potential outcomes Treatment
$i$ $X$ $Y(1)$ $Y(0)$ $T$
1 $x_{1}$ $y_{1}(1)$??? $1$
2 $x_{2}$??? $y_{2}(0)$ 00
3 $x_{3}$??? $y_{3}(0)$ 00
4 $x_{4}$ $y_{4}(1)$??? $1$
5 $x_{5}$ $y_{5}(1)$??? $1$
$\vdots$ $\vdots$ $\vdots$ $\vdots$ $\vdots$
$N$ $x_{N}$??? $y_{N}(0)$ 00
II-A Inferring Counterfactuals through POM
To infer the counterfactuals through POM, one common approach is matching, e.g., exact matching. For a unit $i$, one can search the samples with exactly matching covariates $X$ but treatment $T$ is different, and then estimate the counterfactual outcome $Y_{i}(1-T_{i})$. But it is difficult to conduct if the dimension of $X$ is high, that requires large number of samples to support the matching. Thus, approximate matching is commonly better, e.g., caliper matching [28, 29, 30, 31] and propensity score matching [32, 33, 34, 35]. They approximately search a group of similar samples for the target unit with one or multiple measurements, and finally calculate the counterfactuals.
Another common approach is data-driven imputation, that views the counterfactuals as missing values, and interpolates them by fitting the observational data. To realize it, linear or nonlinear regressions [36, 37, 38, 39] are commonly used to interpolate the missing values from the trained regressive models on the observational data. Different from this, tensor decomposition approaches [40, 41, 42] recover the missing values by decomposing and reconstructing the sparse data tensor. Moreover, deep generative models [10, 11, 12, 13, 14, 15] are also used to model the uncertainty in data, and generate the missing values by randomly sampling. Note that, the generated missing values are not unique, but a group of values conforming to a random distribution. This is different compared to the regressive approaches and tensor decomposition approaches.
II-B Foundational Assumptions of POM
Before using the above approaches on the framework of POM, four assumptions are necessary to be accepted. The first is the stable unit treatment value assumption (SUTVA), defined as
Assumption 1 (SUTVA [43]).
In TABLE I,
  1. 1.
    there is no interference between units, that is, neither $Y_{i}(1)$ nor $Y_{i}(0)$ is affected by what action any other unit received.
  2. 2.
    there is no hidden versions of treatments, that is, no matter how unit $i$ received treatment $T=1$, the outcome that would be observed would be $Y_{i}(1)$, and similarly for treatment $T=0$.
SUTVA guarantees the identical independence between the experimental units, and for unit $i$, the outcome $Y_{i}$ is only up to its treatment $T_{i}$, not the others’. However, this assumption is not always satisfied, e.g., in the case of social network, in which there are frequent interactions between units with varying time.
The second assumption is consistency, defined as
Assumption 2 (Consistency [44, 45]).
In TABLE I, if the $i$-th unit is selected for a treatment $T_{i}$, the observed value of $Y_{i}$, neither $Y_{i}(1)$ nor $Y_{i}(0)$, is the same for all assignments of treatments to the other experimental units.
Another form to describe the consistency assumption through expectation $\mathbb{E}[\cdot]$ is
$\mathbb{E}[Y(1)|T=1]=\sum_{i=1}^{N}y_{i}\times t_{i}=\mathbb{E}[Y|T=1],$ (1)
$\mathbb{E}[Y(0)|T=0]=\sum_{i=1}^{N}y_{i}\times(1-t_{i})=\mathbb{E}[Y|T=0],$
where $y_{1},\dots,y_{N}$ are the samples for $N$ units, and $t_{1},\dots,t_{N}\in\{0,1\}$ are the values of treatment variable $T$, as shown in TABLE I. This is the same to the case of conditional expectation with $X$. Thus, the consistency assumption guarantees the missing counterfactual values in Section II-A can be interpolated by the other observational values actually.
The third assumption is positivity, defined as
Assumption 3 (Positivity [32]).
In TABLE I, $10<P(T=1|X=x)<$ 1.
This means for $X=x$, there are always some treatments randomly assigned with $T=0$ and $T=1$. Otherwise, the counterfactuals for $T=1$ could not be obtained if all units with $X=x$ are assigned with $T=0$, and this is the same to the counterfactuals for $T=0$.
The forth assumption is ignorability, defined as
Assumption 4 (Ignorability [44, 45]).
In TABLE I, $T{\perp\!\!\!\perp}Y(1),Y(0)|X$ for each unit.
Thus, the ignorability assumption implies that all the confounders have been observed in $X$, that is, no hidden confounder would interfere the observations of $Y(1),Y(0)$. Thus, the ignorability assumption is also called as unconfoundedness. This assumption are usually made to avoid the preference on experimental manipulation, that would guarantee the randomization in the experiments [46].
III Structural Causal Model
An SCM is built on a set of variables, including endogenous variables $\mathbf{V}=\{X_{1},\dots,X_{n}\}$ and exogenous variables $\mathbf{U}=\{U_{1},\dots,U_{n}\}$. Note that, endogenous variables are observable from data, but exogenous variables are not. Moreover, a set of functions $\mathbf{F}=\{f_{1},\dots,f_{n}\}$ are also set to describe the functional relationship between these variables.
For example, as shown in Fig. 2, the SCM can be functionally described as
$f_{1}:$ $X_{1}=U_{1},$ (2)
$f_{2}:$ $X_{2}=aX_{1}+U_{2},$
$f_{3}:$ $X_{3}=bX_{1}+cX_{2}+U_{3},$
where $a=0.5,b=0.7,c=0.4$. $U_{1},U_{2},U_{3}$ are all additive noise, and they are mutually independent.
Note that, in an SCM, the functional relationship must conform to a directed-acyclic graphical constraint. The graph is early originated from the concept of Bayesian network [47, 48, 49, 50, 51]. And latter, it is also called as causal graphical model to highlight the causality [52, 53, 54, 16, 24]. And recently, it is also called as causal network in the field of causal discovery [55, 56, 57, 58, 59]. But actually, they are the same in the framework of SCM, and they are all directed acyclic graph (DAG) that represents Markovian knowledge. Thus, to avoid misunderstanding, we here use the name of Bayesian network uniformly, due to the fact that causal graphical model and causal network are not always a Bayesian network in Fig. 2. For example, Direct Cyclic Graph [60, 61, 62, 63], Markov network [64, 65], and Full Time Graph [25], they are also causal graph or causal network.
Figure 2: An example of SCM [24].
Figure 3: A diagram of Pearl’s causal ladder [24].
III-A Pearl’s Causal Ladder [16, 24, 26]
To infer the counterfactuals through SCM, there are three steps:
  1. 1.
    Abduction: Infer the values of exogenous variables $\mathbf{U}$ from the observational data;
  2. 2.
    Action: Perform an intervention on SCM, e.g., $do(X_{2}=2)$ on SCM in Fig. 2, and then, $X_{2}$ would be assigned with value $2$, and the arrows to $X_{2}$ would be modified, as shown in Fig. 3;
  3. 3.
    Prediction: Use the modified SCM to recalculate the counterfactuals of the target variable.
As shown in Fig. 3, we query “observed $X_{1}=0.5,X_{2}=1,X_{3}=1.5$, if had done $X_{2}=2$, what would $X_{3}=?$”. Obviously, this is a counterfactual question. Thus, follow the Pearl’s causal ladder, as shown in Fig. 3, we first infer the values of exogenous variables $\mathbf{U}=\{U_{1},U_{2},U_{3}\}$ from the observational data $X_{1}=0.5,X_{2}=1,X_{3}=1.5$, as follows:
$U_{1}$ $=0.5,$ (3)
$U_{2}$ $=1-0.5\times 0.5=0.75,$
$U_{3}$ $=1.5-0.7\times 0.5-0.4\times 1=0.75.$
Then, $do(X_{2}=2)$ to obtain a modified SCM. And finally, recalculate the counterfactual of $X_{3}$, that is,
$X_{3}(do(X_{2}=2))=0.5\times 0.7+2\times 0.4+0.75=1.9,$ (4)
where $X_{3}(do(X_{2}=2))$ is different compared to the observational $X_{3}$.
The intervention, also named as $do$-operator [16], can be generalized into the form of probability, if modularity assumption is introduced as follow:
Assumption 5 (Modularity [16]).
If a set of variables $\{X_{j_{1}},\dots,X_{j_{p}}\}\subset\mathbf{V}$ is intervened, then for each variable $X\in\mathbf{V}$, it is obtained that
  1. 1.
    if $X\notin\{X_{j_{1}},\dots,X_{j_{p}}\}$, then $P(X|\mathbf{Pa}(X))$ remains unchanged. Here, $\mathbf{Pa}(X)$ is the predecessors of $X$ in Bayesian network, that is also called as causal parents.
  2. 2.
    if $X\in\{X_{j_{1}},\dots,X_{j_{p}}\}$, then $P(X=x|\mathbf{Pa}(X))=1$ if $x$ is the value set by intervention to $X$, otherwise, $P(X=x|\mathbf{Pa}(X))=0$.
This means, if a group of interventions are conducted on $X_{j_{1}},\dots,X_{j_{p}}$, the values of them would be fixed. The connections between them and their causal parents would be broken, and their causal parents would not affect them anymore in the modified SCM, just like in Fig. 3. Moreover, this also means that the probability distributions of the other variables that are not be intervened would not change.
III-B Foundational Assumptions of Bayesian Network
As presented above, to clearly define the causality, every SCM is associate to a Bayesian network with directed-acyclic graphical constraint. Thus, the causality in Bayesian network determines the counterfactuals foundationally. Thus, we discuss the foundational assumptions of Bayesian network in the following contents.
First of all, a concept of d-separation can be defined on a Bayesian network $\mathcal{G}$ as
Definition 1 (d-separation [47, 48, 49, 50, 51]).
Let $\mathbf{X}$, $\mathbf{Y}$ and $\mathbf{Z}$ be three disjoint subsets of endogenous variables $\mathbf{V}=\{X_{1},\dots,X_{n}\}$, and let $\mathbf{p}$ be any path from a node in $\mathbf{X}$ to a node in $\mathbf{Y}$ regardless of direction. $\mathbf{Z}$ is said to block $\mathbf{p}$ if and only if there is a node $v\in\mathbf{p}$ satisfying one of the following items:
  1. 1.
    $v$ has v-structure (two nodes $a,b\in\mathbf{p}$ pointing to $v$, namely $a\rightarrow v\leftarrow b$), and neither $v$ nor its any descendants are in $\mathbf{Z}$;
  2. 2.
    $v$ in $\mathbf{Z}$ and $v$ does not have v-structure.
Then, $\mathbf{Z}$ d-separate $\mathbf{X}$ and $\mathbf{Y}$, denoted as $\mathbf{X}{\perp\!\!\!\perp}_{\mathcal{G}}\mathbf{Y}|\mathbf{Z}$, if $\mathbf{Z}$ blocks any $\mathbf{p}$.
Then, two assumptions, causal Markov (or Markov property) and faithfulness, are introduced as follows:
Assumption 6 (Causal Markov [47, 48, 49, 50, 51]).
Probability distribution $P$ is Markovian to a Bayesian network $\mathcal{G}$ if
$\mathbf{X}{\perp\!\!\!\perp}_{\mathcal{G}}\mathbf{Y}|\mathbf{Z}\Rightarrow\mathbf{X}{\perp\!\!\!\perp}\mathbf{Y}|\mathbf{Z},$ (5)
where $\mathbf{X}$, $\mathbf{Y}$ and $\mathbf{Z}$ are three disjoint subsets.
Assumption 7 (Faithfulness [47, 48, 49, 50, 51]).
Probability distribution $P$ is faithful to a Bayesian network $\mathcal{G}$ if
$\mathbf{X}{\perp\!\!\!\perp}\mathbf{Y}|\mathbf{Z}\Rightarrow\mathbf{X}{\perp\!\!\!\perp}_{\mathcal{G}}\mathbf{Y}|\mathbf{Z}$ (6)
for all disjoint subsets $\mathbf{X}$, $\mathbf{Y}$ and $\mathbf{Z}$.
It is intuitive to understand the two assumptions, that is, the Bayesian network $\mathcal{G}$ has one-to-one correspondence to the independence of probability distribution $P$ in observational data. Thus, if they are satisfied, the joint distribution can be factorized according to the Bayesian network, as follow:
$P(X_{1},\dots,X_{n})=\prod_{i=1}^{n}P(X_{i}|\mathbf{Pa}(X_{i})),$ (7)
where $\mathbf{Pa}(X_{i})$ is the predecessors of $X_{i}$ in $\mathcal{G}$, that is also called as causal parents.
Moreover, a Bayesian network is also assumed to be causally sufficient, that is
Assumption 8 (Causal Sufficiency [47, 48, 49, 50, 51]).
Variables $\mathbf{V}$ are said to satisfy causal sufficiency if there is no hidden variable that is a common cause of two or more variables in $\mathbf{V}$.
This means, if a Bayesian network (or an SCM) is causally sufficient, there is no hidden path connecting two endogenous variables in $\mathbf{V}$, and all variable information is collected sufficiently to support the network. Actually, the ignorability assumption in POM (see Assumption 4) also implies the causal sufficiency, and this assumption is weaker than ignorability. But note that, the causal sufficiency is hard to satisfy practically, because many real-world systems (e.g., economic system and climate system) are complex, thus, we could not collect all the information to describe the systems usually. Thus, some research works focus on this issue. For example, FCI (or Fast Causal Inference) algorithm and its variants [66, 67, 68, 69, 70, 71] are proposed to discover causality in the presence of hidden variables. And some other related works [72, 73, 74] investigate the causal discovery with soft intervention, that intervenes the SCM but does not change the network structure. There are also some related works [75, 76, 77, 78] proposed approaches to search the optimal and efficient adjustment set, that all its variables are observable, minimal and valid, and the removal of any of its variables would destroys the validity.
IV Differences between POM and SCM
Through the above comparisons between POM and SCM, one can find many differences intuitively. For example, they have different statements for counterfactuals, even if some of them can be equally transformed for each other. Moreover, SCM has a causal graph to represent the causations, and the interventions are also introduced to modify the graph to obtain the counterfactuals, but POM does not use it. However, although they have made many distinctions to distinguish each other, as stated in Rubin’s and Pearl’s books [17, 26], there are still some fundamental differences and relationship that need to be discussed.
IV-A Ignorability and Back-door Criterion
In the view of statistics, the ignorability assumption (see Assumption 4) is equivalent in calculation to the causal sufficiency assumption (see Assumption 8) with a back-door criterion. The accurate definition of back-door criterion can be found in these papers and books [53, 54, 16, 24]. Here we present another from with respect to the TABLE I, defined as
Definition 2 (Back-door Criterion [53, 54, 16, 24]).
In TABLE I, the covariate $X$ is said to satisfy back-door criterion, if a Bayesian network is built on all the variables relative to $X,Y,T$, and there is an ordered pair $T\rightarrow Y$ that satisfy
  1. 1.
    no variable in $X$ is a descendant of $T$;
  2. 2.
    $X$ blocks every path between $T$ and $Y$ that contains an arrow into $T$.
Thus, for $X,Y,T$, only one structure can satisfy the back-door criterion, as shown in Fig. 4. And in the back-door path, it is obtained that
$P(Y|do(T=1))$ (8)
$=$ $\sum_{x}P(Y|do(T=1),X=x)P(X=x|do(T=1))$
$=$ $\sum_{x}P(Y|T=1,X=x)P(X=x),$
which is the same in case of $do(T=0)$ [53, 54, 16, 24]. Then, if the causal sufficiency is satisfied, that is, no hidden variable can interfere $X,Y,T$, the expectation of potential outcomes in TABLE I can be calculated by
$\mathbb{E}[Y|do(T=1)]$ (9)
$=$ $\sum_{y}y\times P(Y=y|do(T=1))$
$=$ $\sum_{y}\sum_{x}y\times P(Y=y|T=1,X=x)P(X=x)$
$=$ $\sum_{x}P(X=x)\sum_{y}y\times P(Y=y|T=1,X=x)$
$=$ $\mathbb{E}[\mathbb{E}[Y|T=1,X]].$
This is the same in POM, as follow:
$\mathbb{E}[Y(1)]$ (10)
$=$ $\mathbb{E}[\mathbb{E}[Y(1)|X]]$ $/*Law~{}of~{}full~{}expectation*/$
$=$ $\mathbb{E}[\mathbb{E}[Y(1)|T=1,X]]$ $/*Ignorability*/$
$=$ $\mathbb{E}[\mathbb{E}[Y|T=1,X]],$ $/*Consistency*/$
which is the same in case of $do(T=0)$. Thus, with comparing Eq. (9) and Eq. (10), $\mathbb{E}[Y|do(T=1)]=\mathbb{E}[Y(1)]$, and $\mathbb{E}[Y|do(T=0)]=\mathbb{E}[Y(0)]$, in the two frameworks, but they start from different assumptions. Thus, the ignorability assumption can be decomposed as the causal sufficiency assumption and the back-door criterion, and the causal sufficiency is relatively trivial to satisfy. Actually, the ignorability is more likely to be a technical assumption, that needs many controlled trials to deconfound, while the back-door path is clear to search in a Bayesian network. A potential and more difficult technical issue is to discover the network structure accurately.
IV-B Counterfactual Falsifiability
Back to the example in Fig. 3, if we query to a POM, “Observed $X_{1}=0.5,X_{2}=1,X_{3}=1.5$, if had done $X_{2}=2$, what would $X_{3}=?$”, we must ensure there are at least two observational samples, and they have the same values $X_{1}=0.5$ but different values in $X_{2}$. That is, one is $X_{2}=1$, and another one is $X_{2}=2$, to satisfy the positivity assumption (see Assumption 3). And then, due to the consistency assumption (see Assumption 2), we can infer the counterfactuals through matching the two samples exactly. But this is different in SCM, we can infer the counterfactuals through the Pearl’s causal ladder, even if we only have a single sample, as shown in Section III-A. Thus, how to falsify the counterfactuals?
Figure 4: A diagram of back-door path.
Figure 5: A diagram of data extrapolation.
It is a difficult question, because the counterfactuals are unobserved according to the definition (see “?” in TABLE I), and we need at least two samples for comparison. In general, we cannot back to the past to falsify it with a different treatment, and we cannot reproduce the same experiments completely at different times. Thus, the inferred counterfactuals in SCM cannot be falsified without the assumptions of positivity and consistency, and this view is rejected ambiguously in Pearl’s book [26].
Another explanation is that, without the positivity, but with the consistency only in POM, the counterfactuals can be extrapolated by regressive approaches, and they are falsifiable. For example, as shown in Fig. 5, if there are $T$ and $Y$ only, and the observational data $(T,Y)=\{(0,0.5),(1,1),(2,2.5),(3,2)\}$ is collected, the linear regressive model can be fitted as $Y=0.5T+0.5$. Then, we query “if had done $T=4$, what would $Y=?$”, the counterfactual answer is $Y=2.5$, and the answer can be falsified by the regressive function because it does not beyond the observation. Thus, the consistency assumption actually claims that the counterfactuals can be falsified by the observations. This is also true for SCM, because SCM also uses the regressive approaches to infer counterfactuals.
Moreover, there may be some confusions regarding the preset Bayesian network (see Fig. 2). To build an SCM, a Bayesian network is necessary to be known first. However, how to discover a probabilistic network from a single sample? This is usually impossible. A reasonable explanation is that, the network structure can be discovered from the system mechanisms (e.g., physical mechanisms, communication connections), or graphical knowledge learned from other domains, instead of probability. However, if it is accepted, the definition of causality could be broader and more ambiguous. This is why probabilistic and causal Bayesian network are often indistinguishable, as stated in Pearl’s book [26].
Thus, actually, SCM provides a way to think the counterfactuals, and it emphasizes the thinking way is counterfactual, but not the thinking result, because the result cannot always be falsified without the positivity and consistency assumptions. And on the contrary, POM emphasizes the thinking result is counterfactual. This is a fundamental difference in the problem of counterfactual falsifiability.
IV-C Direct and Indirect Causation
Suppose that there is a causal ordered pathway from one treatment variable to another outcome variable. Then, if there are one or more variables between the two endpoints, the pathway is called as indirect causation, otherwise, it is called as directed causation. Actually, the indirect causation is also called as mediating effect, that has been discussed deeply in Rubin’s papers [79, 80, 81]. Here, Baron-Kenny model [82, 83] is introduced to present mediating analysis for potential outcome with graph. As shown in Fig. 6, $T\rightarrow Y$ is direct causation, and $T\rightarrow X\rightarrow Y$ is indirect causation.
Figure 6: A diagram of mediating analysis [82]. Here, $a,b,c$ are coefficients and $U_{X},U_{Y}$ are biases. $b^{\prime}$ and $U_{total}$ are the coefficient and bias respectively for total effect.
To analyse indirect causation, causal-steps approach [82] is proposed, through building three linear regressive models, as shown in Fig. 6. Suppose that the regressive models are fitted sufficiently, the indirect causation would be detected if satisfy
  1. 1.
    coefficients $a,b$ and $c$ are significant;
  2. 2.
    $|b|<|b^{\prime}|$.
Another approaches are to test the significance of $H_{0}:b^{\prime}-b$ [84, 85, 86]. Or, to test the significance of $H_{0}:ab=0$ [87, 88]. If the correct regressive models are built, the counterfactuals can be calculated, just like SCM in Fig. 2. Moreover, they are all regression-based approaches, that restrict the usage in large multivariate datasets, because one needs to test whether a covariate is a confounder or a mediator, and the test process would be time-consuming seriously.
Figure 7: An example of Forward Counterfactual Inference algorithm. Here, $X_{i}$ is the same as $X_{i}^{\prime}$ to distinguish two recursions.
But on the contrary, the test process can be faster with Bayesian network in an SCM. Here, we provide an algorithm to realize forward counterfactual inference as shown in Algorithm 1, and provide an example to show the inference process as shown in Fig. 7.
The approaches to realize the codes in lines 1 and 3 in Algorithm 1 can be various, e.g., some variants of depth-first-search algorithm and breadth-first-search algorithm [89]. Thus, their time complexities are both $\mathcal{O}(|\mathbf{V}|+|\mathbf{E}|)$, if a Bayesian network $\mathcal{G}=(\mathbf{V},\mathbf{E})$ is given, where $\mathbf{V}$ is the node set and $\mathbf{E}$ is the edge set. Meanwhile, if the computational cost of factorization and adjustment are overlooked, the time complexity of the code in line 9 in Algorithm 1 is $\mathcal{O}(|\mathbf{V}|)$, because we only need to traverse all mediators in each pathway forward, and this process can be performed in parallel. Thus, the total time complexity of Algorithm 1 is $\mathcal{O}(|\mathbf{V}|^{2}+|\mathbf{V}||\mathbf{E}|)$. Thus, if the network structure of the Bayesian network is large, and it can be discovered accurately, the efficiency of SCM to infer counterfactuals would be better with polynomial time complexity.
Input: A Bayesian network $\mathcal{G}$ built on $\mathbf{V}=\{X_{1},\dots,X_{n}\}$ with one target $X_{i}$ and multiple sources $X_{j_{1}},\dots,X_{j_{p}}$.
Output: Distribution $P(X_{i}|do(X_{j_{1}}),\dots,do(X_{j_{p}}))$.
1
2Search one or multiple ordered causal pathways in $\mathcal{G}$ from the sources $X_{j_{1}},\dots,X_{j_{p}}$ to the target $X_{i}$;
3if $X_{j_{1}},\dots,X_{j_{p}}$ are all the direct causation for $X_{i}$ then
4       Search back-door paths (common causes) starting from $X_{j_{1}},\dots,X_{j_{p}}$ and $X_{i}$ reversely in $\mathcal{G}$.
5       if no back-door path within them then
6             return $P(X_{i}|X_{j_{1}},\dots,X_{j_{p}})$;
7      else
8             For every back-door path, select a group of direct predecessors, $\mathbf{W}\subset\mathbf{V}$, and then, according to back-door criterion (see Definition 2),
9             return $\sum_{w}P(X_{i}|X_{j_{1}},\dots,X_{j_{p}},W)P(W)$;
10
11
12Otherwise, for $q$ pathways starting from $X_{j_{p}}$, select $q$ first mediators $X_{j_{p_{1}}},\dots,X_{j_{p_{q}}}$. Then, a group of source-mediator pairs, e.g., $(X_{j_{p}},X_{j_{p_{q}}})$, can be obtained;
13Factorize $P(X_{i}|do(X_{j_{1}}),\dots,do(X_{j_{p}}))$ as
$P(X_{i}|do(X_{j_{1}}),\dots,do(X_{j_{p}}))$ (11)
$=$ $\sum_{x_{j_{1_{1}}:j_{p_{q}}}}P(X_{i}|do(X_{j_{1_{1}}}),\dots,do(X_{j_{2_{1}}}),\dots,do(X_{j_{p_{q}}}))$
×$\times\prod_{p}\prod_{q}P(X_{j_{p_{q}}}|do(X),X\in\mathbf{Pa}(X_{j_{p_{q}}})\subseteq\{X_{j_{1}:j_{p}}\}),$
where $\mathbf{Pa}(X_{j_{p_{q}}})$ is the causal parents of $X_{j_{p_{q}}}$;
14Solve each distribution $P(\cdot|do(\cdot),\dots,do(\cdot))$ in Eq. (11) by Algorithm 1 recursively.
return $P(X_{i}|do(X_{j_{1}}),\dots,do(X_{j_{p}}))$.
Algorithm 1 Forward Counterfactual Inference
V Spatio-Temporal Graphical Counterfactuals
Back to the examples of counterfactual questions at the beginning of this article, that is, if a different investment strategy had been implemented, would we obtain higher returns? And, in computer networks, what changes would occur in network load if a node’s configuration had never been changed? And, would someone still purchase the product even if they had never been shown the advertisement? In these real-world scenarios, interactive behaviors with lagged causal effects are widely-existing. POM cannot answer these questions, because POM does not allow the interactive behaviors between experimental units (see SUTVA in Assumption 1), even with some temporal quasi-experimental approaches (e.g., Regression Discontinuity Design [90, 91] and Differences in Differences [92, 93]). And this is the same to SCM, because the foundational Bayesian network is a DAG that also does not allow mutual connections between network nodes. Thus, to answer these counterfactual queries for interactive units and lagged causal effects, a concept of spatio-temporal graphical counterfactuals is proposed here.
V-A Spatio-Temporal Bayesian Networks
Before the buildings of SCM, actually, various causal graphical models have been proposed to model the spatio-temporal causality. As shown in Fig. 8, Temporal Bayesian Network [64, 94] are proposed to model the momentary causal dependency in a first-order Markov process. Further, Dynamical Bayesian Networks [64] are proposed to represent a combination of multiple Temporal Bayesian Networks. Full Time Graph [25] is proposed to summarize the temporal Bayesian dependency at full timestamps. Note that, Full Time Graph allows the instantaneous causal effects between two units (or variables) in a high-order Markov process, but it only allows strictly stationary causality, that is, the causal dependencies do not change with time. Moreover, Full Time Graph is proved to uniquely decomposed into multiple high-order Temporal Bayesian Networks, named UCN, if the instantaneous causal effects is not allowed [95]. The comparisons of these causal graphs are shown in TABLE II.
Figure 8: Diagrams of various causal graphs.
TABLE II: Comparisons of various causal graphs.
Temporal Bayesian Network Dynamical Bayesian Networks Full Time Graph Spatio-Temporal Bayesian Networks
High-order causation $\times$ $\times$ $\checkmark$ $\checkmark$
Instantaneous causation $\times$ $\checkmark$ $\checkmark$ $\checkmark$
Nonstationary causation $\times$ $\checkmark$ $\times$ $\checkmark$
Thus, based on the characteristics of these causal graphs, Spatio-Temporal Bayesian Networks (STBNs) are defined as a group of DAGs with different timestamps, as shown in Fig. 8. In STBNs, intuitively, the nodes are not equivalent to the variables, but represent the state of the variables at some time steps. With the assumptions of causal Markov (see Assumption 6) and faithfulness (see Assumption 7), the joint distribution of all nodes can be factorized as
$P(X_{1,t},\dots,X_{n,t},X_{1,t-1},\dots,X_{n,t-T+1})$ (12)
$=$ $\prod_{\tau=0}^{T-1}\prod_{i=1}^{n}P(X_{i,t-\tau}|\mathbf{Pa}(X_{i,t-\tau})),$
where $X_{1},\dots,X_{n}$ are the variables, and the timestamps are in range of $10\sim T-$ 1. Theoretically, infinite time steps are allowed in STBNs, but it is usually finite in practice.
To guarantee the global and unique DAG in STBNs, a temporal assumption is introduced as
Assumption 9 (Temporal Assumption).
In STBNs, the cause node precedes or parallels the effect node in time.
This assumption does not allow the directed cause-effect pairs from now to the past, e.g., $X_{i,t-\tau_{1}}\rightarrow X_{j,t-\tau_{2}},\tau_{1}<\tau_{2}$. And this guarantees the functionally identifiable network structure without Markov equivalence class [25, 95, 96]. Moreover, the causal sufficiency assumption (see Assumption 8) is also needed to guarantee the unconfoundedness.
Then, if STBNs are given, the spatial-temporal counterfactuals can be inferred through Algorithm 1, because the global and local STBNs are both DAG. Compared with the naive Bayesian networks, STBNs have the spatio-temporal concept, and the inferring process is naturally forward along with time. This enables to define the momentary interventions for every time steps, and all the interventions would work sequentially, instead of at the same time, like in Algorithm 1. If fully consider this property, the inference algorithm can be faster.
V-B Nonstationarity of STBNs
Nonstationarity is a significant topic in researches on time series. Actually, however, the reasons inducing it are likely to be different from the view of causality. Three of the most currently popular reasons are:
  1. 1.
    Trending: The time series have ascending or descending trending for long terms.
  2. 2.
    Time-varying Variance: The variances of additive noise are time-varying, that is also called as heterogeneous causality in papers [97, 98].
  3. 3.
    Nonstationary Causality: The structures of STBNs change over time, and at least once, as shown in Fig. 9.
Figure 9: Diagrams of nonstationary causality and heterogeneous causality. Here, $C_{t-\tau}=t-\tau,\tau=0,\dots,T-1$ are usually assumed, thus, they are observable, instead of hidden variable.
For the trending time series, an effective approach is to calculate differences of time series in preprocessing. This approach enables to obtain the differential time series, that is the observations of system dynamics, like Eq. (13). This implied, the dynamics is stable during the period of time, and the structures of STBNs do not change, or in another word, the differential time series are stationary. Thus, if the dynamics equations, e.g., Eq. (13), are not equal to zero, the causal links in STBNs are identifiable through the Jacobian matrix of dynamics [99, 100, 101, 102].
For the case of time-varying variance, the causal identifiability is up to the mutually independent exogenous additive noises [96]. That is, if there is no causal connection between two additive noise nodes, the causation is identifiable. Thus, an effective approach is to relate the time-varying noises with a common timestamp, and the timestamps at all time steps are observable confounders $C_{t-\tau},\tau=0,\dots,T-1$ [103, 97, 104, 98]. As shown in Fig. 9, there are three causal pathways, that is, $C_{t-1}\rightarrow U_{4,t-1}\rightarrow X_{4,t-1}$, $C_{t-1}\rightarrow C_{t}\rightarrow U_{3,t}\rightarrow X_{3,t}$, and $C_{t-1}\rightarrow C_{t}\rightarrow U_{4,t}\rightarrow X_{4,t}$. Thus, if $C_{t}$ and $C_{t-1}$ are hidden, the spurious connection between $X_{4,t-1}$ and $X_{3,t}$ would be detected, like the dotted line in Fig. 9. However, $C_{t}$ and $C_{t-1}$ are usually known as the timestamp $t$ and $t-1$, thus, $C_{t}$ and $C_{t-1}$ block the fork paths between $X_{4,t-1},X_{3,t}$ and $X_{4,t}$. Then, the noise variables $U_{4,t-1},U_{3,t}$ and $U_{4,t}$ are mutually independent. Thus, it is causally identifiable. Note that, however, the premise to do so is that any hidden confounders in the case of time-varying variance can be written as smooth or non-smooth functions of timestamps [98].
Nonstationary causality is challenging but critical, especially for the case of frequently changing structures. To identify the nonstationary causality, one approach is to use sliding windows to identify causation within different time periods, as shown in Fig. 9. For each sliding window, it is assumed that the causal structure is invariant, and then, a group of samples would be input to identification algorithms, e.g., [105, 106, 107, 108, 95], to identify a invariant structure during this period. However, if the network changes are frequent, the samples used for identification during a short-term period would be few, but in contrast, we usually need sufficient samples to accept the significance of causal connections. Thus, another reasonable approach is to identify the change points first, and then, identify the causality with the corresponding periods [109, 110]. However, this still does not work in the extreme case of frequently changing, because the samples between two change points may be few. Moreover, there are some other approaches to view the nonstationary causality as a probabilistic normalized flow, and then use variational Bayesian inference to estimate it globally [111, 112, 113].
V-C STBNs and Complex networks
Here, we suppose STBNs model has been built, and it is associated with interactive time-series process of multivariate (or multiple units). Then, we further introduce complex network, another type of network model, to view the causal interactions in complex system. The network dynamics for variables $X_{i},\dots,X_{n}$ can be defined as
$\frac{dX_{i,t}}{dt}=F(X_{i,t})+\sum_{j=1}^{n}A_{ij}G(X_{i,t},X_{j,t}),$ (13)
where $F(\cdot)$ is the self-governed function of $X_{i}$, and $G(\cdot,\cdot)$ is the interactive function for node pairs $X_{i}$ and $X_{j}$. Moreover, $A_{ij}\geq 0$ is the connection between $X_{i}$ and $X_{j}$. Thus, if Eq. (13) is discretized, we can find it the same as first-order STBNs on stationary time series, as shown in Fig. 10.
Thus, the first-order STBNs have some good properties if network dynamics have. For example, if we want to know how to affect a node in STBNs by “intervening” another node, the practical inference steps may not be long, due to the small-world property [114, 115] or the scale-free property [116] of complex networks. Moreover, the counterfactual outcomes would not change because network systems have tolerance from external “interventions” [117, 118]. We would also know the synchronization [119, 120, 121, 122], controllability [123, 124, 125, 126, 127, 128, 129, 130], resilience [131, 132] of STBNs. We would also be able to sparsely identify the network structure even if the network size is large [6, 7], because the independence are equivalent to predictability in STBNs [95]. Also, the network size of STBNs can be reduced, because many complex networks can be simplified and still provide an insightful description of the causality of interest [133, 134]. This is meaningful to design counterfactual inference algorithm on smaller networks, and then obtain higher inference efficiency.
Figure 10: A diagram of discretized network dynamics. Here, $A$ is the adjacency matrix of the first-order STBNs, and $A_{ii}^{\prime},i=1,\dots,n$ represents the joined causal effect of $F$ and $G$.
VI Conclusion
In this work, we mainly focus on the spatio-temporal graphical counterfactuals, and organize an overview for it to discuss its theoretical foundations and application approaches. To discuss the theoretical foundations, a survey is investigated, and the definition of counterfactuals is defined based on the concept of potential outcome in the framework of POM firstly. And then, the framework of SCM is also introduced to infer counterfactuals through graphical languages, which are equivalent to the counterfactuals in POM. Further, to infer counterfactuals on intelligent machine autonomously, a Forward Counterfactual Inference algorithm is designed in this work, and it is able to recursively solve the counterfactual probability distribution, $P(X_{i}|do(X_{j_{1}}),\dots,do(X_{j_{p}}))$, with polynomial time complexity if multi-nodes interventions are conducted on $X_{j_{1}},\dots,X_{j_{p}}$. With this algorithm, the spatio-temporal graphical counterfactuals can be inferred if have two elements, that is, the network structure of STBNs and the corresponding algorithm to identify the network structure. This work not only discusses the identification algorithms in various nonstationary cases, but also discusses the feasibility to improve the algorithm efficiency from the view of complex networks.

### v2

@@
-Spatio-Temporal Graphical Counterfactuals: An Overview
+Kang M, Chen D, Pu Z, Gao J, Yu W. Spatio-Temporal Graphical Counterfactuals: An Overview
 Abstract
@@
 Moreover, there is a lack of graphical approach to infer spatio-temporal counterfactuals, that considers spatial and temporal interactions between multiple units.
-Thus, in this work, our aim is to investigate a survey to compare and discuss different counterfactual models, theories and approaches, and further build a unified graphical causal frameworks to infer the spatio-temporal counterfactuals.
-Index Terms:
-Counterfactual inference, spatio-temporal graphical model, Potential Outcome Model, Structural Causal Model, causality.
-I Introduction
-How to enable an intelligent machine to think and answer a counterfactual question?
+Thus, in this work, our aim is to investigate a survey to compare and discuss different counterfactual models, theories and approaches, and further build a unified graphical causal framework to infer the spatio-temporal counterfactuals.
+keywords:
+Counterfactual inference, spatio-temporal graphical model, Potential Outcome Model, Structural Causal Model, causality
+1 Introduction
+How to enable an intelligent machine to think and answer a spatio-temporal counterfactual question?
 E.g., sometimes we want to know, if a different investment strategy had been implemented, would we obtain higher returns?
@@
 Or, would someone still purchase the product even if they had never been shown the advertisement?
-These questions generally have the counterfactual pattern, that is, “Observed $\dots$, if had done $\dots$, how would $\dots$ go?”.
+These questions generally have the spatio-temporal counterfactual pattern, that is, “Observed $\dots$, if had done $\dots$, how would $\dots$ go?”.
 Thus, according to the pattern, firstly, time is necessary to consider, because the questions are queried now, but the counterfactual actions are supposed to be done in the past.
@@
 But the counterfactual outcomes are not observed, and they are imaged through supposing different actions had been done.
-Figure 1: Infer counterfactuals with less and less knowledge.
-To enable an intelligent machine to imitate counterfactual thinking like human, there are mainly three approaches, as shown in Fig.
+Thirdly, the real-world system can be observed as multivariate time series, and there are widespread multivariate interactions within the time series [1, 2].
+To enable an intelligent machine to imitate counterfactual thinking like human, thus, Potential Outcome Model (POM) [3] and Structural Causal Model (SCM) [4, 5, 6, 7] are proposed as two frameworks with different foundations.
+POM uses stable unit treatment value assumption (SUTVA), consistency assumption, positivity assumption and ignorability assumption as foundations [8, 9, 10, 11].
+If these assumptions are satisfied, one can use matching or imputation approaches to infer counterfactuals without confounders.
+But in contrast, SCM builds a causal ladder, i.e., abduction, action and prediction, to infer counterfactuals through graphical language d-separation [4, 5, 7].
+With the graphical language d-separation, then causal Markov assumption, faithfulness assumption and causal sufficiency assumption are proposed to guarantee the correctness of counterfactual inference on SCM [12, 13, 14, 15, 16].
+However, the different foundations between POM and SCM lead to differences in counterfactual inference.
+Such differences include: (i) POM assumes the positivity, but SCM does not.
+That means the counterfactuals in POM are factual actually and can be found in observations certainly.
+But in contrast, the counterfactuals in SCM can be out of observations and unfalsifiable, that is, one cannot validate the correctness of the inferred counterfactuals.
+(ii) POM assumes the consistency, but it is rejected for SCM [7].
+Actually, SCM also needs the consistency assumption implicitly to guarantee that the counterfactual outcome is one of the unobserved potential outcomes.
+(iii) SCM uses a type of graphical language to deconfound with certain rules, but POM mainly relies on human-guided deconfounding, that makes it difficult to satisfy the ignorability.
+But in contrast, SCM can deconfound through back-door criterion [17, 18, 4, 5] based on the casual sufficiency assumption, and it is equivalent to the ignorability actually.
+(iv) It is difficult for POM to analyze indirect causation (mediation), thus POM typically relies on Baron-Kenny model [19, 20] for assistance.
+By contrast, it is easier for SCM by using front-door criterion [17, 18, 4, 5], that decomposes a causal pathway into two parts, i.e., from source variable to mediator and from mediator to target variable, and then conduct back-door adjustment twice if no other mediator.
+Moreover, whether it is POM or SCM currently, they are both unable to model the spatio-temporal interactions between system units.
+E.g., the SUTVA requires no interaction between system units in POM.
+Moreover, SCM based on the traditional Bayesian network has no modeling for multivariate time series, thus the nonstationarity in time series is often neglected.
+Thus, this motivates us to find a common point to unify the two frameworks, and strengthen the inference for the spatio-temporal counterfactuals.
+Fortunately, our recent work [21, 22] has proposed a Spatio-Temporal Bayesian Networks (STBNs) model based on temporal assumption, that assumes the cause precedes or parallels the effect in time.
+And it is found that, the famous complex networks [23, 24, 25], a type of network dynamics to model spatio-temporal interactions, are a special case of STBNs.
+Thus, we further conduct this work to build a spatio-temporal graphical counterfactual framework based on STBNs.
+The framework requires two components: (i) an STBNs used as the foundation of SCM, (ii) a Forward Counterfactual Inference Algorithm designed on the decomposability of the front-door criterion.
+Then we discuss how to get the two points, following the storyline as shown in Fig.
 1.
-The first is system identification, which aims to identify the system dynamics, and then solve the evolution trajectories with different initial values [1].
-This process may be data-driven and adaptive, but the physical mechanisms and models are necessary to be known as knowledge in prior [2, 3, 4, 5].
-E.g., in [6, 7], functional basis or network structure are necessary to be known at least one to identify the network dynamics sparsely.
-However, the knowledge of system mechanism is usually hard to obtain in prior.
-The second approach is generative modeling, which aims to model and learn the system uncertainty, and then generate scenarios with different random inputs [8, 9].
-Thus, compared with system identification, generative approach is not necessary to know the determined part, that is the knowledge of system mechanism, but on contrary, it focuses on the uncertain modeling, especially with the deep learning [10, 11, 12, 13, 14, 15].
-However, the generated scenarios would not be out of the distribution of the observational data [9], while the counterfactual scenarios are usually not in the observational factuals.
-Thus, the mechanism knowledge is still needed to correct the bias between the distributions of factuals and counterfactuals.
-Thus, the third approach is causal inference, which aims to discover causality from the observational data, and then infer factuals and counterfactuals directly with as little knowledge as possible [16, 17].
-Causal counterfactual inference enables artificial intelligence to learn the knowledge from data, and ultimately improve their performances for new scenarios.
-To realize it, many research works have been proposed to infer counterfactuals, including the temporal cases [18, 19, 20, 21, 22, 23].
-However, there are still two challenging but critical issues to address: (i) Most current inference approaches are built on the frameworks of Potential Outcome Model (POM) [17] or Structural Causal Model (SCM) [16, 24, 25, 26].
-The goals of them are the same, but their foundations are different.
-This induces difficulty in discussing them in a unified framework.
-(ii) The two frameworks both cannot infer spatio-temporal counterfactuals, that considers spatial and temporal interactions between system units.
 Thus, the main contributions of this work are as follows:
   1. 1.
-    A survey is investigated to discuss the frameworks from POM to SCM, including their theoretical foundations and application approaches. This part is organized from Section II to Section IV.
+    A survey is investigated to discuss the frameworks from POM to SCM, including their theoretical foundations and application approaches. This part is organized from Section 2 to Section 4.
   2. 2.
     A Forward Counterfactual Inference algorithm is designed here to recursively and autonomously infer counterfactuals through causal graphical languages.
-This part is presented as Algorithm 1 in Section IV-C.
+This part is presented as Algorithm 1 in Section 4.3.
   3. 3.
     An overview of the spatio-temporal graphical counterfactual framework is proposed to discuss Spatio-Temporal Bayesian Networks, its nonstationarity, and the relationship with complex networks.
-This part is organized in Section V.
-II Potential Outcome Model
-The counterfactuals are defined as the unobserved potential outcomes in POM, as shown in TABLE I.
+This part is organized in Section 5.
+Figure 1: Towards spatio-temporal graphical counterfactuals.
+2 Potential Outcome Model
+The counterfactuals are defined as the unobserved potential outcomes in POM, as shown in Table 1.
 Here, $Y$ is the potential outcomes of the target variable.
@@
 In this case, for different treatments $T$, only one potential outcome, $Y(1)$ or $Y(0)$, can be observed, and another one is unobserved, namely the counterfactual.
-TABLE I: Data for POM example.
+Table 1: Data for POM example.
    Units    Covariates    Potential outcomes    Treatment
    $i$    $X$    $Y(1)$    $Y(0)$    $T$
-1 $x_{1}$ $y_{1}(1)$??? $1$
-2 $x_{2}$??? $y_{2}(0)$ 00
-3 $x_{3}$??? $y_{3}(0)$ 00
-4 $x_{4}$ $y_{4}(1)$??? $1$
-5 $x_{5}$ $y_{5}(1)$??? $1$
+   1    $x_{1}$    $y_{1}(1)$    ??    11
+   2    $x_{2}$    ??    $y_{2}(0)$    0
+   3    $x_{3}$    ??    $y_{3}(0)$    0
+   4    $x_{4}$    $y_{4}(1)$    ??    11
+   5    $x_{5}$    $y_{5}(1)$    ??    11
    $\vdots$    $\vdots$    $\vdots$    $\vdots$    $\vdots$
-$N$ $x_{N}$??? $y_{N}(0)$ 00
-II-A Inferring Counterfactuals through POM
+   NN    $x_{N}$    ??    $y_{N}(0)$    0
+2.1 Inferring Counterfactuals through POM
 To infer the counterfactuals through POM, one common approach is matching, e.g., exact matching.
@@
 Different from this, tensor decomposition approaches [38, 39, 40] recover the missing values by decomposing and reconstructing the sparse data tensor.
-Moreover, deep generative models [10, 11, 12, 13, 14, 15] are also used to model the uncertainty in data, and generate the missing values by randomly sampling.
+Moreover, deep generative models [41, 42, 43, 44, 45, 46, 47] are also used to model the uncertainty in data, and generate the missing values by randomly sampling.
 Note that, the generated missing values are not unique, but a group of values conforming to a random distribution.
 This is different compared to the regressive approaches and tensor decomposition approaches.
-II-B Foundational Assumptions of POM
-Before using the above approaches on the framework of POM, four assumptions are necessary to be accepted. The first is the stable unit treatment value assumption (SUTVA), defined as
-Assumption 1 (SUTVA [43]).
-In TABLE I,
+2.2 Foundational Assumptions of POM
+Before using the above approaches on the framework of POM, four assumptions are necessary to be accepted. The first is SUTVA, which is defined as {assumption}[SUTVA [10]] In Table 1,
   1. 1.
@@
 However, this assumption is not always satisfied, e.g., in the case of social network, in which there are frequent interactions between units with varying time.
-The second assumption is consistency, defined as
-Assumption 2 (Consistency [44, 45]).
-In TABLE I, if the $i$-th unit is selected for a treatment $T_{i}$, the observed value of $Y_{i}$, neither $Y_{i}(1)$ nor $Y_{i}(0)$, is the same for all assignments of treatments to the other experimental units.
+The second assumption is consistency, defined as {assumption}[Consistency [8, 9]] In Table 1, if the $i$-th unit is selected for a treatment $T_{i}$, the observed value of $Y_{i}$, neither $Y_{i}(1)$ nor $Y_{i}(0)$, is the same for all assignments of treatments to the other experimental units.
 Another form to describe the consistency assumption through expectation $\mathbb{E}[\cdot]$ is
@@
 $\mathbb{E}[Y(0)|T=0]=\sum_{i=1}^{N}y_{i}\times(1-t_{i})=\mathbb{E}[Y|T=0],$
-where $y_{1},\dots,y_{N}$ are the samples for $N$ units, and $t_{1},\dots,t_{N}\in\{0,1\}$ are the values of treatment variable $T$, as shown in TABLE I.
+where $y_{1},\dots,y_{N}$ are the samples for $N$ units, and $t_{1},\dots,t_{N}\in\{0,1\}$ are the values of treatment variable $T$, as shown in Table 1.
 This is the same to the case of conditional expectation with $X$.
-Thus, the consistency assumption guarantees the missing counterfactual values in Section II-A can be interpolated by the other observational values actually.
-The third assumption is positivity, defined as
-Assumption 3 (Positivity [32]).
-In TABLE I, $10<P(T=1|X=x)<$ 1.
+Thus, the consistency assumption guarantees the missing counterfactual values in Section 2.1 can be interpolated by the other observational values actually.
+The third assumption is positivity, defined as {assumption}[Positivity [11]] In Table 1, $0<P(T=1|X=x)<1$.
 This means for $X=x$, there are always some treatments randomly assigned with $T=0$ and $T=1$.
 Otherwise, the counterfactuals for $T=1$ could not be obtained if all units with $X=x$ are assigned with $T=0$, and this is the same to the counterfactuals for $T=0$.
-The forth assumption is ignorability, defined as
-Assumption 4 (Ignorability [44, 45]).
-In TABLE I, $T{\perp\!\!\!\perp}Y(1),Y(0)|X$ for each unit.
+The forth assumption is ignorability, defined as {assumption}[Ignorability [8, 9]] In Table 1, $T{\perp\!\!\!\perp}Y(1),Y(0)|X$ for each unit.
 Thus, the ignorability assumption implies that all the confounders have been observed in $X$, that is, no hidden confounder would interfere the observations of $Y(1),Y(0)$.
@@
 This assumption are usually made to avoid the preference on experimental manipulation, that would guarantee the randomization in the experiments [48].
-III Structural Causal Model
+3 Structural Causal Model
+Different from POM, SCM provides a type of graphical language to discuss causality intuitively, which avoids complex human-guided technical operations.
 An SCM is built on a set of variables, including endogenous variables $\mathbf{V}=\{X_{1},\dots,X_{n}\}$ and exogenous variables $\mathbf{U}=\{U_{1},\dots,U_{n}\}$.
@@
 Figure 2: An example of SCM [5].
-Figure 3: A diagram of Pearl’s causal ladder [24].
-III-A Pearl’s Causal Ladder [16, 24, 26]
+3.1 Pearl’s Causal Ladder [4, 5, 7]
 To infer the counterfactuals through SCM, there are three steps:
@@
   2. 2.
-    Action: Perform an intervention on SCM, e.g., $do(X_{2}=2)$ on SCM in Fig. 2, and then, $X_{2}$ would be assigned with value $2$, and the arrows to $X_{2}$ would be modified, as shown in Fig. 3;
+    Action: Perform an intervention on SCM, e.g., $do(X_{2}=2)$ on SCM in Fig. 2, and then, $X_{2}$ would be assigned with value 22, and the arrows to $X_{2}$ would be modified, as shown in Fig. 3;
   3. 3.
     Prediction: Use the modified SCM to recalculate the counterfactuals of the target variable.
+Figure 3: A diagram of Pearl’s causal ladder [5].
 As shown in Fig.
@@
 3, we first infer the values of exogenous variables $\mathbf{U}=\{U_{1},U_{2},U_{3}\}$ from the observational data $X_{1}=0.5,X_{2}=1,X_{3}=1.5$, as follows:
-$U_{1}$ $=0.5,$ (3)
-$U_{2}$ $=1-0.5\times 0.5=0.75,$
-$U_{3}$ $=1.5-0.7\times 0.5-0.4\times 1=0.75.$
+[…]
+[…]
+[…]
 Then, $do(X_{2}=2)$ to obtain a modified SCM. And finally, recalculate the counterfactual of $X_{3}$, that is,
-$X_{3}(do(X_{2}=2))=0.5\times 0.7+2\times 0.4+0.75=1.9,$ (4)
+[…]
 where $X_{3}(do(X_{2}=2))$ is different compared to the observational $X_{3}$.
-The intervention, also named as $do$-operator [16], can be generalized into the form of probability, if modularity assumption is introduced as follow:
-Assumption 5 (Modularity [16]).
-If a set of variables $\{X_{j_{1}},\dots,X_{j_{p}}\}\subset\mathbf{V}$ is intervened, then for each variable $X\in\mathbf{V}$, it is obtained that
+The intervention, also named as $do$-operator [4], can be generalized into the form of probability, if modularity assumption is introduced as follow: {assumption}[Modularity [4]] If a set of variables $\{X_{j_{1}},\dots,X_{j_{p}}\}\subset\mathbf{V}$ is intervened, then for each variable $X\in\mathbf{V}$, it is obtained that
   1. 1.
@@
 Moreover, this also means that the probability distributions of the other variables that are not be intervened would not change.
-III-B Foundational Assumptions of Bayesian Network
+3.2 Foundational Assumptions of Bayesian Network
 As presented above, to clearly define the causality, every SCM is associate to a Bayesian network with directed-acyclic graphical constraint.
@@
 First of all, a concept of d-separation can be defined on a Bayesian network $\mathcal{G}$ as
-Definition 1 (d-separation [47, 48, 49, 50, 51]).
+Definition 3.1 (d-separation [12, 13, 14, 15, 16]).
 Let $\mathbf{X}$, $\mathbf{Y}$ and $\mathbf{Z}$ be three disjoint subsets of endogenous variables $\mathbf{V}=\{X_{1},\dots,X_{n}\}$, and let $\mathbf{p}$ be any path from a node in $\mathbf{X}$ to a node in $\mathbf{Y}$ regardless of direction.
@@
 Then, $\mathbf{Z}$ d-separate $\mathbf{X}$ and $\mathbf{Y}$, denoted as $\mathbf{X}{\perp\!\!\!\perp}_{\mathcal{G}}\mathbf{Y}|\mathbf{Z}$, if $\mathbf{Z}$ blocks any $\mathbf{p}$.
-Then, two assumptions, causal Markov (or Markov property) and faithfulness, are introduced as follows:
-Assumption 6 (Causal Markov [47, 48, 49, 50, 51]).
-Probability distribution $P$ is Markovian to a Bayesian network $\mathcal{G}$ if
+Then, two assumptions, causal Markov (or Markov property) and faithfulness, are introduced as follows: {assumption}[Causal Markov [12, 13, 14, 15, 16]] Probability distribution $P$ is Markovian to a Bayesian network $\mathcal{G}$ if
 $\mathbf{X}{\perp\!\!\!\perp}_{\mathcal{G}}\mathbf{Y}|\mathbf{Z}\Rightarrow\mathbf{X}{\perp\!\!\!\perp}\mathbf{Y}|\mathbf{Z},$ (5)
 where $\mathbf{X}$, $\mathbf{Y}$ and $\mathbf{Z}$ are three disjoint subsets.
-Assumption 7 (Faithfulness [47, 48, 49, 50, 51]).
-Probability distribution $P$ is faithful to a Bayesian network $\mathcal{G}$ if
+{assumption}[Faithfulness [12, 13, 14, 15, 16]] Probability distribution $P$ is faithful to a Bayesian network $\mathcal{G}$ if
 $\mathbf{X}{\perp\!\!\!\perp}\mathbf{Y}|\mathbf{Z}\Rightarrow\mathbf{X}{\perp\!\!\!\perp}_{\mathcal{G}}\mathbf{Y}|\mathbf{Z}$ (6)
@@
 where $\mathbf{Pa}(X_{i})$ is the predecessors of $X_{i}$ in $\mathcal{G}$, that is also called as causal parents.
-Moreover, a Bayesian network is also assumed to be causally sufficient, that is
-Assumption 8 (Causal Sufficiency [47, 48, 49, 50, 51]).
-Variables $\mathbf{V}$ are said to satisfy causal sufficiency if there is no hidden variable that is a common cause of two or more variables in $\mathbf{V}$.
+Moreover, a Bayesian network is also assumed to be causally sufficient, that is {assumption}[Causal Sufficiency [12, 13, 14, 15, 16]] Variables $\mathbf{V}$ are said to satisfy causal sufficiency if there is no hidden variable that is a common cause of two or more variables in $\mathbf{V}$.
 This means, if a Bayesian network (or an SCM) is causally sufficient, there is no hidden path connecting two endogenous variables in $\mathbf{V}$, and all variable information is collected sufficiently to support the network.
-Actually, the ignorability assumption in POM (see Assumption 4) also implies the causal sufficiency, and this assumption is weaker than ignorability.
+Actually, the ignorability assumption in POM (see Assumption 2.2) also implies the causal sufficiency, and this assumption is weaker than ignorability.
 But note that, the causal sufficiency is hard to satisfy practically, because many real-world systems (e.g., economic system and climate system) are complex, thus, we could not collect all the information to describe the systems usually.
@@
 There are also some related works [70, 71, 72, 73] proposed approaches to search the optimal and efficient adjustment set, that all its variables are observable, minimal and valid, and the removal of any of its variables would destroys the validity.
-IV Differences between POM and SCM
-Through the above comparisons between POM and SCM, one can find many differences intuitively.
-For example, they have different statements for counterfactuals, even if some of them can be equally transformed for each other.
-Moreover, SCM has a causal graph to represent the causations, and the interventions are also introduced to modify the graph to obtain the counterfactuals, but POM does not use it.
-However, although they have made many distinctions to distinguish each other, as stated in Rubin’s and Pearl’s books [17, 26], there are still some fundamental differences and relationship that need to be discussed.
-IV-A Ignorability and Back-door Criterion
-In the view of statistics, the ignorability assumption (see Assumption 4) is equivalent in calculation to the causal sufficiency assumption (see Assumption 8) with a back-door criterion.
+4 Differences between POM and SCM
+In this section, the fundamental differences between POM and SCM are discussed, including (i) counterfactual falsifiability: POM assumes the positivity, thus the counterfactuals in POM can be falsified on observations, but SCM does not.
+The counterfactuals in SCM can be outside the observations, thus it is unfalsifiable.
+(ii) using back-door criterion to satisfy ignorability: POM requires ignorability to guarantee the unconfoundedness and correctness of counterfactual inference, but it is a complex human-guided operation.
+By contrast, SCM realizes unconfoundedness by the causal sufficiency and back-door adjustment, that are equivalent to the ignorability assumption.
+SCM facilitates counterfactual inference through graphical language.
+(iii) direct causation and indirect causation: POM struggles to process the indirect causation, and it typically relies on Baron-Kenny model.
+But in contrast, SCM has advantages to adjust the front-door path (that has indirect causation) through equivalently applying the back-door criterion twice.
+We further use this advantage to design the forward counterfactual inference algorithm to infer the spatio-temporal counterfactuals autonomously.
+The following are the details.
+4.1 Counterfactual Falsifiability
+As shown in Fig.
+3, if we query to a POM, “Observed $X_{1}=0.5,X_{2}=1,X_{3}=1.5$, if had done $X_{2}=2$, what would $X_{3}=?$”, we must ensure there are at least two observational samples, and they have the same values $X_{1}=0.5$ but different values in $X_{2}$.
+That is, one is $X_{2}=1$, and another one is $X_{2}=2$, to satisfy the positivity assumption (see Assumption 2.2).
+And then, due to the consistency assumption (see Assumption 2.2), we can infer the counterfactuals through matching the two samples exactly.
+But this is different in SCM, we can infer the counterfactuals through the Pearl’s causal ladder, even if we only have a single sample, as shown in Section 3.1.
+Thus, how to falsify the counterfactuals?
+It is a difficult question, because the counterfactuals are unobserved according to the definition (see “?” in Table 1), and we need at least two samples for comparison.
+In general, we cannot back to the past to falsify it with a different treatment, and we cannot reproduce the same experiments completely at different times.
+Thus, the inferred counterfactuals in SCM cannot be falsified without the assumptions of positivity and consistency, and this view is rejected ambiguously in Pearl’s book [7].
+Another explanation is that, without the positivity, but with the consistency only in POM, the counterfactuals can be extrapolated by regressive approaches, and they are falsifiable.
+For example, as shown in Fig.
+5, if there are $T$ and $Y$ only, and the observational data $(T,Y)=\{(0,0.5),(1,1),(2,2.5),(3,2)\}$ is collected, the linear regressive model can be fitted as $Y=0.5T+0.5$.
+Then, we query “if had done $T=4$, what would Y=?Y=?”, the counterfactual answer is $Y=2.5$, and the answer can be falsified by the regressive function because it does not beyond the observation.
+Thus, the consistency assumption actually claims that the counterfactuals can be falsified by the observations.
+This is also true for SCM, because SCM also uses the regressive approaches to infer counterfactuals.
+Moreover, there may be some confusions regarding the preset Bayesian network (see Fig.
+2).
+To build an SCM, a Bayesian network is necessary to be known first.
+However, how to discover a probabilistic network from a single sample?
+This is usually impossible.
+A reasonable explanation is that, the network structure can be discovered from the system mechanisms (e.g., physical mechanisms, communication connections), or graphical knowledge learned from other domains, instead of probability.
+However, if it is accepted, the definition of causality could be broader and more ambiguous.
+This is why probabilistic and causal Bayesian network are often indistinguishable, as stated in Pearl’s book [7].
+Thus, actually, SCM provides a way to think the counterfactuals, and it emphasizes the thinking way is counterfactual, but not the thinking result, because the result cannot always be falsified without the positivity and consistency assumptions.
+And on the contrary, POM emphasizes the thinking result is counterfactual.
+This is a fundamental difference in the problem of counterfactual falsifiability.
+4.2 Ignorability and Back-door Criterion
+In the view of statistics, the ignorability assumption (see Assumption 2.2) is equivalent in calculation to the causal sufficiency assumption (see Assumption 3.1) with a back-door criterion.
 The accurate definition of back-door criterion can be found in these papers and books [17, 18, 4, 5].
-Here we present another from with respect to the TABLE I, defined as
-Definition 2 (Back-door Criterion [53, 54, 16, 24]).
-In TABLE I, the covariate $X$ is said to satisfy back-door criterion, if a Bayesian network is built on all the variables relative to $X,Y,T$, and there is an ordered pair $T\rightarrow Y$ that satisfy
+Here we present another from with respect to the Table 1, defined as
+Definition 4.1 (Back-door Criterion [17, 18, 4, 5]).
+In Table 1, the covariate $X$ is said to satisfy back-door criterion, if a Bayesian network is built on all the variables relative to $X,Y,T$, and there is an ordered pair $T\rightarrow Y$ that satisfy
   1. 1.
@@
     $X$ blocks every path between $T$ and $Y$ that contains an arrow into $T$.
+Figure 4: A diagram of data extrapolation.
+Figure 5: A diagram of back-door path.
 Thus, for $X,Y,T$, only one structure can satisfy the back-door criterion, as shown in Fig. 5. And in the back-door path, it is obtained that
@@
 which is the same in case of $do(T=0)$ [17, 18, 4, 5].
-Then, if the causal sufficiency is satisfied, that is, no hidden variable can interfere $X,Y,T$, the expectation of potential outcomes in TABLE I can be calculated by
+Then, if the causal sufficiency is satisfied, that is, no hidden variable can interfere $X,Y,T$, the expectation of potential outcomes in Table 1 can be calculated by
 $\mathbb{E}[Y|do(T=1)]$ (9)
@@
 A potential and more difficult technical issue is to discover the network structure accurately.
-IV-B Counterfactual Falsifiability
-Back to the example in Fig.
-3, if we query to a POM, “Observed $X_{1}=0.5,X_{2}=1,X_{3}=1.5$, if had done $X_{2}=2$, what would $X_{3}=?$”, we must ensure there are at least two observational samples, and they have the same values $X_{1}=0.5$ but different values in $X_{2}$.
-That is, one is $X_{2}=1$, and another one is $X_{2}=2$, to satisfy the positivity assumption (see Assumption 3).
-And then, due to the consistency assumption (see Assumption 2), we can infer the counterfactuals through matching the two samples exactly.
-But this is different in SCM, we can infer the counterfactuals through the Pearl’s causal ladder, even if we only have a single sample, as shown in Section III-A.
-Thus, how to falsify the counterfactuals?
-Figure 4: A diagram of back-door path.
-Figure 5: A diagram of data extrapolation.
-It is a difficult question, because the counterfactuals are unobserved according to the definition (see “?” in TABLE I), and we need at least two samples for comparison.
-In general, we cannot back to the past to falsify it with a different treatment, and we cannot reproduce the same experiments completely at different times.
-Thus, the inferred counterfactuals in SCM cannot be falsified without the assumptions of positivity and consistency, and this view is rejected ambiguously in Pearl’s book [26].
-Another explanation is that, without the positivity, but with the consistency only in POM, the counterfactuals can be extrapolated by regressive approaches, and they are falsifiable.
-For example, as shown in Fig.
-5, if there are $T$ and $Y$ only, and the observational data $(T,Y)=\{(0,0.5),(1,1),(2,2.5),(3,2)\}$ is collected, the linear regressive model can be fitted as $Y=0.5T+0.5$.
-Then, we query “if had done $T=4$, what would $Y=?$”, the counterfactual answer is $Y=2.5$, and the answer can be falsified by the regressive function because it does not beyond the observation.
-Thus, the consistency assumption actually claims that the counterfactuals can be falsified by the observations.
-This is also true for SCM, because SCM also uses the regressive approaches to infer counterfactuals.
-Moreover, there may be some confusions regarding the preset Bayesian network (see Fig.
-2).
-To build an SCM, a Bayesian network is necessary to be known first.
-However, how to discover a probabilistic network from a single sample?
-This is usually impossible.
-A reasonable explanation is that, the network structure can be discovered from the system mechanisms (e.g., physical mechanisms, communication connections), or graphical knowledge learned from other domains, instead of probability.
-However, if it is accepted, the definition of causality could be broader and more ambiguous.
-This is why probabilistic and causal Bayesian network are often indistinguishable, as stated in Pearl’s book [26].
-Thus, actually, SCM provides a way to think the counterfactuals, and it emphasizes the thinking way is counterfactual, but not the thinking result, because the result cannot always be falsified without the positivity and consistency assumptions.
-And on the contrary, POM emphasizes the thinking result is counterfactual.
-This is a fundamental difference in the problem of counterfactual falsifiability.
-IV-C Direct and Indirect Causation
+4.3 Direct and Indirect Causation
 Suppose that there is a causal ordered pathway from one treatment variable to another outcome variable.
@@
 As shown in Fig.
-6, $T\rightarrow Y$ is direct causation, and $T\rightarrow X\rightarrow Y$ is indirect causation.
+7, $T\rightarrow Y$ is direct causation, and $T\rightarrow X\rightarrow Y$ is indirect causation.
 Figure 6: A diagram of mediating analysis [19]. Here, $a,b,c$ are coefficients and $U_{X},U_{Y}$ are biases. $b^{\prime}$ and $U_{total}$ are the coefficient and bias respectively for total effect.
+Figure 7: A diagram of front-door adjustment [5]. Here, $T$ is a treatment variable, and $Y$ is a target variable. $X_{1}$ and $X_{2}$ are both covariates.
 To analyse indirect causation, causal-steps approach [19] is proposed, through building three linear regressive models, as shown in Fig.
-6.
+7.
 Suppose that the regressive models are fitted sufficiently, the indirect causation would be detected if satisfy
@@
     $|b|<|b^{\prime}|$.
-Another approaches are to test the significance of $H_{0}:b^{\prime}-b$ [84, 85, 86].
+Another approaches are to test the significance of $H_{0}:b^{\prime}-b=0$ [77, 78, 79].
 Or, to test the significance of $H_{0}:ab=0$ [80, 81].
@@
 Moreover, they are all regression-based approaches, that restrict the usage in large multivariate datasets, because one needs to test whether a covariate is a confounder or a mediator, and the test process would be time-consuming seriously.
-Figure 7: An example of Forward Counterfactual Inference algorithm. Here, $X_{i}$ is the same as $X_{i}^{\prime}$ to distinguish two recursions.
 But on the contrary, the test process can be faster with Bayesian network in an SCM.
-Here, we provide an algorithm to realize forward counterfactual inference as shown in Algorithm 1, and provide an example to show the inference process as shown in Fig.
-7.
-The approaches to realize the codes in lines 1 and 3 in Algorithm 1 can be various, e.g., some variants of depth-first-search algorithm and breadth-first-search algorithm [89].
+Here, we propose an algorithm to realize forward counterfactual inference as shown in Algorithm 1.
+To clearly demonstrate the principle of Algorithm 1, we first introduce the front-door criterion as follows:
+Definition 4.2 (Front-door Criterion [17, 18, 4, 5]).
+The covariate $X$ is said to satisfy front-door criterion, if a Bayesian network is built on all the variables relative to $X,Y,T$, and there is an ordered pair $T\rightarrow Y$ that satisfy
+  1. 1.
+    $X$ intercepts all directed paths from $T$ to $Y$;
+  2. 2.
+    there is no back-door path from $T$ to $X$;
+  3. 3.
+    all back-door paths from $X$ to $Y$ are blocked by $T$.
+Then, a classic example is provided to show the front-door adjustment through the theorem 3.4.1 in book [5], as also shown in Fig.
+7 here.
+If given a treatment variable $T$ and a target variable $Y$, with two covariates $X_{1}$ and $X_{2}$.
+Then one can adjust the distribution $P(Y|do(T))$ with the front-door criterion as follows:
+$P(Y|do(T))$ $=\sum_{x_{1}}P(X_{1}|do(T))P(Y|do(X_{1}))$ (11)
+$=\sum_{x_{1}}P(X_{1}|T)\sum_{t^{\prime}}P(Y|X_{1},T^{\prime})P(T^{\prime})$
+Thus, we can conclude from this equation:
+  1. 1.
+    If there are one or multiple ordered causal pathways from $T$ to $Y$ and there is at least one mediator on the ways, then the causal pathway can be decomposed into two causal pathways through that mediator, e.g., $T\rightarrow X_{1}$ and $X_{1}\rightarrow Y$ in Eq.
+(11).
+  2. 2.
+    If there is no mediator on all causal pathways, then adjustment using back-door criterion. Of course, if there is no back-door path, no need to adjust.
+Thus, the front-door principle is non-atomic and decomposable.
+And with this more foundational principle, we propose the forward counterfactual algorithm to autonomously apply this adjustment principle, as presented in Algorithm 1.
+The algorithm can be conducted on arbitrary variable set $\{X_{1},\dots,X_{n}\}$ that satisfied causal sufficiency.
+And we also provide an example to show the inference process as shown in Fig.
+8.
+In Algorithm 1, the approaches to realize the codes in lines 1 and 3 can be various, e.g., some variants of depth-first-search algorithm and breadth-first-search algorithm [82].
 Thus, their time complexities are both $\mathcal{O}(|\mathbf{V}|+|\mathbf{E}|)$, if a Bayesian network $\mathcal{G}=(\mathbf{V},\mathbf{E})$ is given, where $\mathbf{V}$ is the node set and $\mathbf{E}$ is the edge set.
@@
 Thus, if the network structure of the Bayesian network is large, and it can be discovered accurately, the efficiency of SCM to infer counterfactuals would be better with polynomial time complexity.
+5 Spatio-Temporal Graphical Counterfactuals
+Then for the examples of counterfactual questions at the beginning of this article, that is, if a different investment strategy had been implemented, would we obtain higher returns?
+And, in computer networks, what changes would occur in network load if a node’s configuration had never been changed?
+And, would someone still purchase the product even if they had never been shown the advertisement?
+In these real-world scenarios, interactive behaviors with lagged causal effects are widely-existing.
+POM cannot answer these questions, because POM does not allow the interactive behaviors between experimental units (see SUTVA in Assumption 2.2), even with some temporal quasi-experimental approaches (e.g., Regression Discontinuity Design [83, 84] and Differences in Differences [85, 86]).
+And this is the same to SCM, because the foundational Bayesian network is a DAG that also does not allow mutual connections between network nodes.
+Thus, to answer these counterfactual queries for interactive units and lagged causal effects, a concept of spatio-temporal graphical counterfactuals is proposed here.
 Input: A Bayesian network $\mathcal{G}$ built on $\mathbf{V}=\{X_{1},\dots,X_{n}\}$ with one target $X_{i}$ and multiple sources $X_{j_{1}},\dots,X_{j_{p}}$.
@@
 7 else
-8             For every back-door path, select a group of direct predecessors, $\mathbf{W}\subset\mathbf{V}$, and then, according to back-door criterion (see Definition 2),
+8      For every back-door path, select a group of direct predecessors, $\mathbf{W}\subset\mathbf{V}$, and then, according to back-door criterion (see Definition 4.1),
 9    return $\sum_{w}P(X_{i}|X_{j_{1}},\dots,X_{j_{p}},W)P(W)$;
@@
 $=$ $\sum_{x_{j_{1_{1}}:j_{p_{q}}}}P(X_{i}|do(X_{j_{1_{1}}}),\dots,do(X_{j_{2_{1}}}),\dots,do(X_{j_{p_{q}}}))$
-×$\times\prod_{p}\prod_{q}P(X_{j_{p_{q}}}|do(X),X\in\mathbf{Pa}(X_{j_{p_{q}}})\subseteq\{X_{j_{1}:j_{p}}\}),$
+$\times\prod_{p}\prod_{q}P(X_{j_{p_{q}}}|do(X),X\in\mathbf{Pa}(X_{j_{p_{q}}})\subseteq\{X_{j_{1}:j_{p}}\}),$
 where $\mathbf{Pa}(X_{j_{p_{q}}})$ is the causal parents of $X_{j_{p_{q}}}$;
@@
 Algorithm 1 Forward Counterfactual Inference
-V Spatio-Temporal Graphical Counterfactuals
-Back to the examples of counterfactual questions at the beginning of this article, that is, if a different investment strategy had been implemented, would we obtain higher returns?
-And, in computer networks, what changes would occur in network load if a node’s configuration had never been changed?
-And, would someone still purchase the product even if they had never been shown the advertisement?
-In these real-world scenarios, interactive behaviors with lagged causal effects are widely-existing.
-POM cannot answer these questions, because POM does not allow the interactive behaviors between experimental units (see SUTVA in Assumption 1), even with some temporal quasi-experimental approaches (e.g., Regression Discontinuity Design [90, 91] and Differences in Differences [92, 93]).
-And this is the same to SCM, because the foundational Bayesian network is a DAG that also does not allow mutual connections between network nodes.
-Thus, to answer these counterfactual queries for interactive units and lagged causal effects, a concept of spatio-temporal graphical counterfactuals is proposed here.
-V-A Spatio-Temporal Bayesian Networks
+Figure 8: An example of Forward Counterfactual Inference algorithm. Here, $X_{i}$ is the same as $X_{i}^{\prime}$ to distinguish two recursions.
+Table 2: Comparisons of various causal graphs.
+Temporal Bayesian Network Dynamical Bayesian Networks Full Time Graph Spatio-Temporal Bayesian Networks
+High-order causation $\times$ $\times$ $\checkmark$ $\checkmark$
+Instantaneous causation $\times$ $\checkmark$ $\checkmark$ $\checkmark$
+Nonstationary causation $\times$ $\checkmark$ $\times$ $\checkmark$
+5.1 Spatio-Temporal Bayesian Networks
 Before the buildings of SCM, actually, various causal graphical models have been proposed to model the spatio-temporal causality.
 As shown in Fig.
-8, Temporal Bayesian Network [64, 94] are proposed to model the momentary causal dependency in a first-order Markov process.
+9, Temporal Bayesian Network [59, 87] are proposed to model the momentary causal dependency in a first-order Markov process.
 Further, Dynamical Bayesian Networks [59] are proposed to represent a combination of multiple Temporal Bayesian Networks.
@@
 Moreover, Full Time Graph is proved to uniquely decomposed into multiple high-order Temporal Bayesian Networks, named UCN, if the instantaneous causal effects is not allowed [21].
-The comparisons of these causal graphs are shown in TABLE II.
+The comparisons of these causal graphs are shown in Table 2.
 Figure 9: Diagrams of various causal graphs.
-TABLE II: Comparisons of various causal graphs.
-Temporal Bayesian Network Dynamical Bayesian Networks Full Time Graph Spatio-Temporal Bayesian Networks
-High-order causation $\times$ $\times$ $\checkmark$ $\checkmark$
-Instantaneous causation $\times$ $\checkmark$ $\checkmark$ $\checkmark$
-Nonstationary causation $\times$ $\checkmark$ $\times$ $\checkmark$
-Thus, based on the characteristics of these causal graphs, Spatio-Temporal Bayesian Networks (STBNs) are defined as a group of DAGs with different timestamps, as shown in Fig.
-8.
+Thus, based on the characteristics of these causal graphs, STBNs are defined as a group of DAGs with different timestamps, as shown in Fig.
+9.
 In STBNs, intuitively, the nodes are not equivalent to the variables, but represent the state of the variables at some time steps.
@@
 $=$ $\prod_{\tau=0}^{T-1}\prod_{i=1}^{n}P(X_{i,t-\tau}|\mathbf{Pa}(X_{i,t-\tau})),$
-where $X_{1},\dots,X_{n}$ are the variables, and the timestamps are in range of $10\sim T-$ 1. Theoretically, infinite time steps are allowed in STBNs, but it is usually finite in practice.
-To guarantee the global and unique DAG in STBNs, a temporal assumption is introduced as
-Assumption 9 (Temporal Assumption).
-In STBNs, the cause node precedes or parallels the effect node in time.
+where $X_{1},\dots,X_{n}$ are the variables, and the timestamps are in range of $0\sim T-1$. Theoretically, infinite time steps are allowed in STBNs, but it is usually finite in practice.
+To guarantee the global and unique DAG in STBNs, a temporal assumption is introduced as {assumption}[Temporal Assumption] In STBNs, the cause node precedes or parallels the effect node in time.
 This assumption does not allow the directed cause-effect pairs from now to the past, e.g., $X_{i,t-\tau_{1}}\rightarrow X_{j,t-\tau_{2}},\tau_{1}<\tau_{2}$.
 And this guarantees the functionally identifiable network structure without Markov equivalence class [6, 21, 88].
-Moreover, the causal sufficiency assumption (see Assumption 8) is also needed to guarantee the unconfoundedness.
+Moreover, the causal sufficiency assumption (see Assumption 3.1) is also needed to guarantee the unconfoundedness.
 Then, if STBNs are given, the spatial-temporal counterfactuals can be inferred through Algorithm 1, because the global and local STBNs are both DAG.
@@
 If fully consider this property, the inference algorithm can be faster.
-V-B Nonstationarity of STBNs
+5.2 Nonstationarity of STBNs
 Nonstationarity is a significant topic in researches on time series.
@@
 As shown in Fig.
-9, there are three causal pathways, that is, $C_{t-1}\rightarrow U_{4,t-1}\rightarrow X_{4,t-1}$, $C_{t-1}\rightarrow C_{t}\rightarrow U_{3,t}\rightarrow X_{3,t}$, and $C_{t-1}\rightarrow C_{t}\rightarrow U_{4,t}\rightarrow X_{4,t}$.
+10, there are three causal pathways, that is, $C_{t-1}\rightarrow U_{4,t-1}\rightarrow X_{4,t-1}$, $C_{t-1}\rightarrow C_{t}\rightarrow U_{3,t}\rightarrow X_{3,t}$, and $C_{t-1}\rightarrow C_{t}\rightarrow U_{4,t}\rightarrow X_{4,t}$.
 Thus, if $C_{t}$ and $C_{t-1}$ are hidden, the spurious connection between $X_{4,t-1}$ and $X_{3,t}$ would be detected, like the dotted line in Fig.
-9.
-However, $C_{t}$ and $C_{t-1}$ are usually known as the timestamp $t$ and $t-1$, thus, $C_{t}$ and $C_{t-1}$ block the fork paths between $X_{4,t-1},X_{3,t}$ and $X_{4,t}$.
+10.
+However, $C_{t}$ and $C_{t-1}$ are usually known as the timestamp $t$ and t1t-1, thus, $C_{t}$ and $C_{t-1}$ block the fork paths between $X_{4,t-1},X_{3,t}$ and $X_{4,t}$.
 Then, the noise variables $U_{4,t-1},U_{3,t}$ and $U_{4,t}$ are mutually independent.
@@
 To identify the nonstationary causality, one approach is to use sliding windows to identify causation within different time periods, as shown in Fig.
-9.
-For each sliding window, it is assumed that the causal structure is invariant, and then, a group of samples would be input to identification algorithms, e.g., [105, 106, 107, 108, 95], to identify a invariant structure during this period.
+10.
+For each sliding window, it is assumed that the causal structure is invariant, and then, a group of samples would be input to identification algorithms, e.g., [97, 98, 99, 100, 21, 22], to identify a invariant structure during this period.
 However, if the network changes are frequent, the samples used for identification during a short-term period would be few, but in contrast, we usually need sufficient samples to accept the significance of causal connections.
@@
 Moreover, there are some other approaches to view the nonstationary causality as a probabilistic normalized flow, and then use variational Bayesian inference to estimate it globally [103, 104, 105].
-V-C STBNs and Complex networks
+5.3 STBNs and Complex networks
 Here, we suppose STBNs model has been built, and it is associated with interactive time-series process of multivariate (or multiple units).
@@
 (14) is discretized, we can find it the same as first-order STBNs on stationary time series, as shown in Fig.
-10.
+11.
+Figure 11: A diagram of discretized network dynamics. Here, $A$ is the adjacency matrix of the first-order STBNs, and $A_{ii}^{\prime},i=1,\dots,n$ represents the joined causal effect of $F$ and $G$.
 Thus, the first-order STBNs have some good properties if network dynamics have.
@@
 This is meaningful to design counterfactual inference algorithm on smaller networks, and then obtain higher inference efficiency.
-Figure 10: A diagram of discretized network dynamics. Here, $A$ is the adjacency matrix of the first-order STBNs, and $A_{ii}^{\prime},i=1,\dots,n$ represents the joined causal effect of $F$ and $G$.
-VI Conclusion
+6 Conclusion
 In this work, we mainly focus on the spatio-temporal graphical counterfactuals, and organize an overview for it to discuss its theoretical foundations and application approaches.
@@
 This work not only discusses the identification algorithms in various nonstationary cases, but also discusses the feasibility to improve the algorithm efficiency from the view of complex networks.
+[…]
+This work is supported by the National Key R&D Program of China under Grant No.
+2022ZD0120003, the National Natural Science Foundation of China under Grant Nos.
+62233004, 62273090, 52525204, 52572354, the Jiangsu Provincial Scientific Research Center of Applied Mathematics under Grant No.
+BK20233002, the US National Science Foundation under Grant No.
+2047488 and the Rensselaer-IBM AI Research Collaboration.

### v3

@@
-Kang M, Chen D, Pu Z, Gao J, Yu W. Spatio-Temporal Graphical Counterfactuals: An Overview
+Kang M Y, Chen D X, Pu Z Y, et al. Spatio-Temporal Graphical Counterfactuals: An Overview
 Abstract
-Counterfactual thinking is a critical yet challenging topic for artificial intelligence to learn knowledge from data and ultimately improve their performances for new scenarios.
-Many research works, including Potential Outcome Model and Structural Causal Model, have been proposed to realize it.
-However, their modelings, theoretical foundations and application approaches are usually different.
-Moreover, there is a lack of graphical approach to infer spatio-temporal counterfactuals, that considers spatial and temporal interactions between multiple units.
-Thus, in this work, our aim is to investigate a survey to compare and discuss different counterfactual models, theories and approaches, and further build a unified graphical causal framework to infer the spatio-temporal counterfactuals.
+Counterfactual thinking is a crucial yet challenging topic for artificial intelligence to learn knowledge from data and ultimately improve performance for new scenarios.
+Many research works, including the Potential Outcome Model (POM) and the Structural Causal Model (SCM), have been proposed to address this.
+However, their modeling, theoretical foundations, and application approaches often differ.
+Moreover, there is a lack of graphical approaches for inferring spatio-temporal counterfactuals, that account for spatial and temporal interactions among multiple units.
+Thus, in this work, we aim to present a survey that compares and discusses different counterfactual models, theories and approaches.
+Additionally, we propose a unified graphical causal framework to infer spatio-temporal counterfactuals.
 keywords:
@@
 1 Introduction
-How to enable an intelligent machine to think and answer a spatio-temporal counterfactual question?
-E.g., sometimes we want to know, if a different investment strategy had been implemented, would we obtain higher returns?
+How can we enable an intelligent machine to think about and answer a spatio-temporal counterfactual question?
+For example, sometimes we want to know: if a different investment strategy had been implemented, would we have obtained higher returns?
 Or, in computer networks, what changes would occur in network load if a node’s configuration had never been changed?
-Or, would someone still purchase the product even if they had never been shown the advertisement?
-These questions generally have the spatio-temporal counterfactual pattern, that is, “Observed $\dots$, if had done $\dots$, how would $\dots$ go?”.
-Thus, according to the pattern, firstly, time is necessary to consider, because the questions are queried now, but the counterfactual actions are supposed to be done in the past.
-Secondly, counterfactual outcomes are unobservable.
-The outcomes of some actions are observed if the actions have been done to the real-world system, thus these outcomes are factual.
-But the counterfactual outcomes are not observed, and they are imaged through supposing different actions had been done.
-Thirdly, the real-world system can be observed as multivariate time series, and there are widespread multivariate interactions within the time series [1, 2].
-To enable an intelligent machine to imitate counterfactual thinking like human, thus, Potential Outcome Model (POM) [3] and Structural Causal Model (SCM) [4, 5, 6, 7] are proposed as two frameworks with different foundations.
+Or would someone still purchase a product even if they had never seen the advertisement?
+These questions generally follow a spatio-temporal counterfactual pattern: “Given the observed $\dots$, if $\dots$ had been done, how would $\dots$ have evolved?”.
+Thus, according to the pattern, time must be considered first, because the questions are posed now, but the counterfactual actions are supposed to be performed in the past.
+Second, counterfactual outcomes are unobservable.
+The outcomes of some actions are observed when they are performed on a real-world system, thus, these outcomes are factual.
+But the counterfactual outcomes are not observed, and they are imagined by supposing that different actions had been taken.
+Third, the real-world system can be observed as a multivariate time series, and there are widespread multivariate interactions within it [1, 2].
+To enable intelligent machines to mimic counterfactual thinking, the Potential Outcome Model (POM) [3] and Structural Causal Model (SCM) [4, 5, 6, 7] have been proposed as two frameworks with different foundations.
 POM uses stable unit treatment value assumption (SUTVA), consistency assumption, positivity assumption and ignorability assumption as foundations [8, 9, 10, 11].
-If these assumptions are satisfied, one can use matching or imputation approaches to infer counterfactuals without confounders.
-But in contrast, SCM builds a causal ladder, i.e., abduction, action and prediction, to infer counterfactuals through graphical language d-separation [4, 5, 7].
-With the graphical language d-separation, then causal Markov assumption, faithfulness assumption and causal sufficiency assumption are proposed to guarantee the correctness of counterfactual inference on SCM [12, 13, 14, 15, 16].
-However, the different foundations between POM and SCM lead to differences in counterfactual inference.
-Such differences include: (i) POM assumes the positivity, but SCM does not.
-That means the counterfactuals in POM are factual actually and can be found in observations certainly.
-But in contrast, the counterfactuals in SCM can be out of observations and unfalsifiable, that is, one cannot validate the correctness of the inferred counterfactuals.
-(ii) POM assumes the consistency, but it is rejected for SCM [7].
-Actually, SCM also needs the consistency assumption implicitly to guarantee that the counterfactual outcome is one of the unobserved potential outcomes.
-(iii) SCM uses a type of graphical language to deconfound with certain rules, but POM mainly relies on human-guided deconfounding, that makes it difficult to satisfy the ignorability.
-But in contrast, SCM can deconfound through back-door criterion [17, 18, 4, 5] based on the casual sufficiency assumption, and it is equivalent to the ignorability actually.
-(iv) It is difficult for POM to analyze indirect causation (mediation), thus POM typically relies on Baron-Kenny model [19, 20] for assistance.
-By contrast, it is easier for SCM by using front-door criterion [17, 18, 4, 5], that decomposes a causal pathway into two parts, i.e., from source variable to mediator and from mediator to target variable, and then conduct back-door adjustment twice if no other mediator.
-Moreover, whether it is POM or SCM currently, they are both unable to model the spatio-temporal interactions between system units.
-E.g., the SUTVA requires no interaction between system units in POM.
-Moreover, SCM based on the traditional Bayesian network has no modeling for multivariate time series, thus the nonstationarity in time series is often neglected.
-Thus, this motivates us to find a common point to unify the two frameworks, and strengthen the inference for the spatio-temporal counterfactuals.
-Fortunately, our recent work [21, 22] has proposed a Spatio-Temporal Bayesian Networks (STBNs) model based on temporal assumption, that assumes the cause precedes or parallels the effect in time.
-And it is found that, the famous complex networks [23, 24, 25], a type of network dynamics to model spatio-temporal interactions, are a special case of STBNs.
+If these assumptions are hold, one can use matching or imputation approaches to infer counterfactuals without confounders.
+In contrast, SCM builds a causal ladder, i.e., abduction, action and prediction, to infer counterfactuals through graphical language d-separation [4, 5, 7].
+With the graphical language d-separation, the causal Markov assumption, faithfulness assumption and causal sufficiency assumption are proposed to guarantee the correctness of counterfactual inference on SCM [12, 13, 14, 15, 16].
+However, the distinct foundations between POM and SCM lead to differences in counterfactual inference.
+These differences include: (i) POM assumes positivity, whereas SCM does not.
+Thus, the counterfactuals in POM are factual and can be found in observations certainly.
+In contrast, counterfactuals in SCM can be out of observations and are therefore unfalsifiable, that is, their correctness cannot be empirically validated.
+(ii) POM explicitly assumes consistency, but this assumption is rejected for SCM [7].
+Actually, SCM also implicitly relies on the consistency assumption to ensure that the counterfactual outcome corresponds to one of the possibly unobserved potential outcomes.
+(iii) SCM uses a type of graphical language to deconfound with certain rules, but POM mainly relies on human-guided deconfounding, which makes it difficult to satisfy the ignorability.
+In contrast, SCM can deconfound via the back-door criterion [17, 18, 4, 5] under the casual sufficiency assumption, and it is equivalent to ignorability.
+(iv) POM faces limitations to analyze indirect causation (mediation), thus POM typically relies on the Baron-Kenny model [19, 20] for assistance.
+By contrast, SCM facilitates this analysis through the front-door criterion [17, 18, 4, 5], which decomposes a causal pathway into two segments: from the source variable to mediator, and from the mediator to target variable.
+Back-door adjustment is then applied twice, assuming no additional mediators are presented.
+Moreover, both POM and SCM currently lack the ability to model the spatio-temporal interactions among system units.
+For example, the SUTVA requires no interaction between system units in POM.
+Similarly, SCM, when based on the traditional Bayesian network, does not handle multivariate time series and often overlooks nonstationarity.
+This limitation motivates us to identify a common point to unify the two frameworks and strengthen inference for spatio-temporal counterfactuals.
+Our recent work [21, 22] has proposed a spatio-temporal Bayesian networks (STBNs) model based on a temporal assumption that the cause precedes or parallels the effect in time.
+It is found that the famous complex networks [23, 24, 25], a type of network dynamics used to model spatio-temporal interactions, are also a special case of STBNs.
 Thus, we further conduct this work to build a spatio-temporal graphical counterfactual framework based on STBNs.
-The framework requires two components: (i) an STBNs used as the foundation of SCM, (ii) a Forward Counterfactual Inference Algorithm designed on the decomposability of the front-door criterion.
-Then we discuss how to get the two points, following the storyline as shown in Fig.
+This framework requires two components: (i) STBNs used as the foundation of SCM, (ii) a Forward Counterfactual Inference Algorithm designed on the decomposability of the front-door criterion.
+We then discuss how to get the two points, following the storyline as shown in Fig.
 1.
@@
   1. 1.
-    A survey is investigated to discuss the frameworks from POM to SCM, including their theoretical foundations and application approaches. This part is organized from Section 2 to Section 4.
+    A survey is conducted to discuss the frameworks from POM to SCM, including their theoretical foundations and application approaches. This part is organized from Section 2 to Section 4.
   2. 2.
@@
   3. 3.
-    An overview of the spatio-temporal graphical counterfactual framework is proposed to discuss Spatio-Temporal Bayesian Networks, its nonstationarity, and the relationship with complex networks.
+    An overview of the spatio-temporal graphical counterfactual framework is proposed to discuss the spatio-temporal Bayesian networks, their nonstationarity, and the relationship with complex networks.
 This part is organized in Section 5.
@@
 The counterfactuals are defined as the unobserved potential outcomes in POM, as shown in Table 1.
-Here, $Y$ is the potential outcomes of the target variable.
-$T$ is treatment variable.
-$T=1$ represents being treated and $T=0$ represents being controlled, with a group of randomized control trials on $N$ units.
+Let $Y$ be the potential outcomes of the target variable, and $T$ is the treatment variable.
+$T=1$ represents assignment to the treatment group, and $T=0$ represents assignment to the control group in a randomized controlled trial involving $N$ units.
 $X$ represents covariates that have relationship with $Y$ and $T$, and are controlled for randomized trials.
@@
 2.1 Inferring Counterfactuals through POM
-To infer the counterfactuals through POM, one common approach is matching, e.g., exact matching.
+To infer the counterfactuals using POM, one common approach is exact sample matching.
 For a unit $i$, one can search the samples with exactly matching covariates $X$ but treatment $T$ is different, and then estimate the counterfactual outcome $Y_{i}(1-T_{i})$.
-But it is difficult to conduct if the dimension of $X$ is high, that requires large number of samples to support the matching.
+But it is difficult to conduct when the dimension of $X$ is high, as it requires a large number of samples to support the matching.
 Thus, approximate matching is commonly better, e.g., caliper matching [27, 28, 29, 30] and propensity score matching [11, 31, 32, 33].
-They approximately search a group of similar samples for the target unit with one or multiple measurements, and finally calculate the counterfactuals.
-Another common approach is data-driven imputation, that views the counterfactuals as missing values, and interpolates them by fitting the observational data.
-To realize it, linear or nonlinear regressions [34, 35, 36, 37] are commonly used to interpolate the missing values from the trained regressive models on the observational data.
-Different from this, tensor decomposition approaches [38, 39, 40] recover the missing values by decomposing and reconstructing the sparse data tensor.
-Moreover, deep generative models [41, 42, 43, 44, 45, 46, 47] are also used to model the uncertainty in data, and generate the missing values by randomly sampling.
-Note that, the generated missing values are not unique, but a group of values conforming to a random distribution.
-This is different compared to the regressive approaches and tensor decomposition approaches.
+They search for a group of similar samples to the target unit with one or more measurements, then calculate counterfactuals.
+Another common approach is data-driven imputation, which views counterfactuals as missing values, and interpolates them by fitting the observational data.
+Linear or nonlinear regressions [34, 35, 36, 37] are typically used to interpolate missing values based on patterns learned from the observational data.
+In contrast, tensor decomposition approaches [38, 39, 40] recover the missing values by decomposing and reconstructing a sparse data tensor.
+Deep generative models [41, 42, 43, 44, 45, 46, 47] have also been used to capture data uncertainty and generate missing values by randomly sampling.
+Note that, the imputed missing values produced by generative models are not unique.
+Instead, they constitute a distribution of plausible values, which distinguishes them from regression-based and tensor decomposition approaches.
 2.2 Foundational Assumptions of POM
-Before using the above approaches on the framework of POM, four assumptions are necessary to be accepted. The first is SUTVA, which is defined as {assumption}[SUTVA [10]] In Table 1,
+Before using the above approaches within POM framework, four assumptions must be accepted. The first is SUTVA, which is defined as {assumption}[SUTVA [10]] In Table 1,
   1. 1.
@@
   2. 2.
-    there is no hidden versions of treatments, that is, no matter how unit $i$ received treatment $T=1$, the outcome that would be observed would be $Y_{i}(1)$, and similarly for treatment $T=0$.
+    There are no hidden versions of treatments, that is, no matter how unit $i$ received treatment $T=1$, the outcome that would be observed would be $Y_{i}(1)$, and similarly for treatment $T=0$.
 SUTVA guarantees the identical independence between the experimental units, and for unit $i$, the outcome $Y_{i}$ is only up to its treatment $T_{i}$, not the others’.
-However, this assumption is not always satisfied, e.g., in the case of social network, in which there are frequent interactions between units with varying time.
+However, this assumption is not always satisfied, e.g., in social networks, where interactions occur at varying times.
 The second assumption is consistency, defined as {assumption}[Consistency [8, 9]] In Table 1, if the $i$-th unit is selected for a treatment $T_{i}$, the observed value of $Y_{i}$, neither $Y_{i}(1)$ nor $Y_{i}(0)$, is the same for all assignments of treatments to the other experimental units.
@@
 where $y_{1},\dots,y_{N}$ are the samples for $N$ units, and $t_{1},\dots,t_{N}\in\{0,1\}$ are the values of treatment variable $T$, as shown in Table 1.
-This is the same to the case of conditional expectation with $X$.
-Thus, the consistency assumption guarantees the missing counterfactual values in Section 2.1 can be interpolated by the other observational values actually.
+This is the same as the case of taking a conditional expectation with respect to $X$.
+Thus, the consistency assumption guarantees that the missing counterfactual values in Section 2.1 can be interpolated by the other observational values.
 The third assumption is positivity, defined as {assumption}[Positivity [11]] In Table 1, $0<P(T=1|X=x)<1$.
-This means for $X=x$, there are always some treatments randomly assigned with $T=0$ and $T=1$.
-Otherwise, the counterfactuals for $T=1$ could not be obtained if all units with $X=x$ are assigned with $T=0$, and this is the same to the counterfactuals for $T=0$.
-The forth assumption is ignorability, defined as {assumption}[Ignorability [8, 9]] In Table 1, $T{\perp\!\!\!\perp}Y(1),Y(0)|X$ for each unit.
-Thus, the ignorability assumption implies that all the confounders have been observed in $X$, that is, no hidden confounder would interfere the observations of $Y(1),Y(0)$.
-Thus, the ignorability assumption is also called as unconfoundedness.
-This assumption are usually made to avoid the preference on experimental manipulation, that would guarantee the randomization in the experiments [48].
+This means that, for $X=x$, some treatments are always assigned to $T=0$ or $T=1$ at random.
+Otherwise, the counterfactuals for $T=1$ could not be obtained if all units with $X=x$ are assigned with $T=0$, and this is the same as the counterfactuals for $T=0$.
+The fourth assumption is ignorability, defined as {assumption}[Ignorability [8, 9]] In Table 1, $T{\perp\!\!\!\perp}Y(1),Y(0)|X$ for each unit.
+Thus, the ignorability assumption implies that all the confounders have been observed in $X$, that is, no hidden confounder would interfere with the observations of $Y(1),Y(0)$.
+Thus, the ignorability assumption is also called unconfoundedness assumption.
+This assumption is usually made to avoid the preference for experimental manipulation, which would guarantee randomization in the experiments [48].
 3 Structural Causal Model
-Different from POM, SCM provides a type of graphical language to discuss causality intuitively, which avoids complex human-guided technical operations.
+In contrast to POM, SCM provides a graphical language for intuitively discussing causality, thereby avoiding complex human-guided technical operations.
 An SCM is built on a set of variables, including endogenous variables $\mathbf{V}=\{X_{1},\dots,X_{n}\}$ and exogenous variables $\mathbf{U}=\{U_{1},\dots,U_{n}\}$.
 Note that, endogenous variables are observable from data, but exogenous variables are not.
-Moreover, a set of functions $\mathbf{F}=\{f_{1},\dots,f_{n}\}$ are also set to describe the functional relationship between these variables.
+Moreover, a set of functions $\mathbf{F}=\{f_{1},\dots,f_{n}\}$ can also be used to describe the functional relationship between these variables.
+Figure 2: Example of SCM [5].
+Figure 3: Diagram of Pearl’s causal ladder [5].
 For example, as shown in Fig. 3, the SCM can be functionally described as
@@
 where $a=0.5,b=0.7,c=0.4$. $U_{1},U_{2},U_{3}$ are all additive noise, and they are mutually independent.
-Note that, in an SCM, the functional relationship must conform to a directed-acyclic graphical constraint.
-The graph is early originated from the concept of Bayesian network [12, 13, 14, 15, 16].
-And latter, it is also called as causal graphical model to highlight the causality [49, 17, 18, 4, 5].
+In an SCM, the functional relationship must conform to the directed acyclic graphical constraint.
+The graph is originally based on the concept of Bayesian network [12, 13, 14, 15, 16].
+And later, it is also called as causal graphical model to highlight its causal structure [49, 17, 18, 4, 5].
 And recently, it is also called as causal network in the field of causal discovery [50, 51, 52, 53, 54].
@@
 Thus, to avoid misunderstanding, we here use the name of Bayesian network uniformly, due to the fact that causal graphical model and causal network are not always a Bayesian network in Fig.
-2.
-For example, Direct Cyclic Graph [55, 56, 57, 58], Markov network [59, 60], and Full Time Graph [6], they are also causal graph or causal network.
-Figure 2: An example of SCM [5].
+3.
+For example, direct cyclic graphs [55, 56, 57, 58], Markov networks [59, 60], and full time graph [6], they are also causal graph or causal network.
 3.1 Pearl’s Causal Ladder [4, 5, 7]
@@
     Prediction: Use the modified SCM to recalculate the counterfactuals of the target variable.
-Figure 3: A diagram of Pearl’s causal ladder [5].
 As shown in Fig.
-3, we query “observed $X_{1}=0.5,X_{2}=1,X_{3}=1.5$, if had done $X_{2}=2$, what would $X_{3}=?$”.
+3, consider the following query “observed $X_{1}=0.5,X_{2}=1,X_{3}=1.5$, if had done $X_{2}=2$, what would $X_{3}=?$”.
 Obviously, this is a counterfactual question.
-Thus, follow the Pearl’s causal ladder, as shown in Fig.
-3, we first infer the values of exogenous variables $\mathbf{U}=\{U_{1},U_{2},U_{3}\}$ from the observational data $X_{1}=0.5,X_{2}=1,X_{3}=1.5$, as follows:
+Thus, following Pearl’s causal ladder, as shown in Fig.
+3, we first infer the values of exogenous variables $\mathbf{U}=\{U_{1},U_{2},U_{3}\}$ from the observational data $X_{1}=0.5,X_{2}=1$ and $X_{3}=1.5$, as follows:
 […]
@@
 where $X_{3}(do(X_{2}=2))$ is different compared to the observational $X_{3}$.
-The intervention, also named as $do$-operator [4], can be generalized into the form of probability, if modularity assumption is introduced as follow: {assumption}[Modularity [4]] If a set of variables $\{X_{j_{1}},\dots,X_{j_{p}}\}\subset\mathbf{V}$ is intervened, then for each variable $X\in\mathbf{V}$, it is obtained that
+The intervention, also named as $do$-operator [4], can be generalized into the form of probability, if the modularity assumption is introduced as follows: {assumption}[Modularity [4]] If a set of variables $\{X_{j_{1}},\dots,X_{j_{p}}\}\subset\mathbf{V}$ is intervened, then for each variable $X\in\mathbf{V}$, it is obtained that
   1. 1.
     if $X\notin\{X_{j_{1}},\dots,X_{j_{p}}\}$, then $P(X|\mathbf{Pa}(X))$ remains unchanged.
-Here, $\mathbf{Pa}(X)$ is the predecessors of $X$ in Bayesian network, that is also called as causal parents.
+Here, $\mathbf{Pa}(X)$ denotes the set of predecessors of $X$ in a Bayesian network, also called the causal parents of $X$.
   2. 2.
     if $X\in\{X_{j_{1}},\dots,X_{j_{p}}\}$, then $P(X=x|\mathbf{Pa}(X))=1$ if $x$ is the value set by intervention to $X$, otherwise, $P(X=x|\mathbf{Pa}(X))=0$.
-This means, if a group of interventions are conducted on $X_{j_{1}},\dots,X_{j_{p}}$, the values of them would be fixed.
+This means that if a set of interventions is applied to $X_{j_{1}},\dots,X_{j_{p}}$, their values would be fixed.
 The connections between them and their causal parents would be broken, and their causal parents would not affect them anymore in the modified SCM, just like in Fig.
 3.
-Moreover, this also means that the probability distributions of the other variables that are not be intervened would not change.
+Moreover, this also means that the probability distributions of the other variables that are not intervened upon would not change.
 3.2 Foundational Assumptions of Bayesian Network
-As presented above, to clearly define the causality, every SCM is associate to a Bayesian network with directed-acyclic graphical constraint.
-Thus, the causality in Bayesian network determines the counterfactuals foundationally.
-Thus, we discuss the foundational assumptions of Bayesian network in the following contents.
+As presented above, to clearly define causality, each SCM is associated with a Bayesian network satisfying the directed acyclic graphical constraint.
+Thus, the causality in the Bayesian network provides the foundation for counterfactuals.
+Thus, we discuss the foundational assumptions of Bayesian networks in the following sections.
 First of all, a concept of d-separation can be defined on a Bayesian network $\mathcal{G}$ as
@@
 for all disjoint subsets $\mathbf{X}$, $\mathbf{Y}$ and $\mathbf{Z}$.
-It is intuitive to understand the two assumptions, that is, the Bayesian network $\mathcal{G}$ has one-to-one correspondence to the independence of probability distribution $P$ in observational data.
-Thus, if they are satisfied, the joint distribution can be factorized according to the Bayesian network, as follow:
+It is intuitive to understand the two assumptions: the Bayesian network $\mathcal{G}$ corresponds one-to-one to the independence structure of the probability distribution $P$ in the observational data.
+Thus, if they are satisfied, the joint distribution can be factorized according to the Bayesian network, as follows:
 $P(X_{1},\dots,X_{n})=\prod_{i=1}^{n}P(X_{i}|\mathbf{Pa}(X_{i})),$ (7)
@@
 Moreover, a Bayesian network is also assumed to be causally sufficient, that is {assumption}[Causal Sufficiency [12, 13, 14, 15, 16]] Variables $\mathbf{V}$ are said to satisfy causal sufficiency if there is no hidden variable that is a common cause of two or more variables in $\mathbf{V}$.
-This means, if a Bayesian network (or an SCM) is causally sufficient, there is no hidden path connecting two endogenous variables in $\mathbf{V}$, and all variable information is collected sufficiently to support the network.
+This means that if a Bayesian network (or an SCM) is causally sufficient, there is no hidden path connecting two endogenous variables in $\mathbf{V}$, and all variable information is collected sufficiently to support the network.
 Actually, the ignorability assumption in POM (see Assumption 2.2) also implies the causal sufficiency, and this assumption is weaker than ignorability.
-But note that, the causal sufficiency is hard to satisfy practically, because many real-world systems (e.g., economic system and climate system) are complex, thus, we could not collect all the information to describe the systems usually.
+But note that, causal sufficiency is hard to satisfy in practice, because many real-world systems (e.g., the economic and climate systems) are complex, thus, we usually cannot collect all the information needed to describe them.
 Thus, some research works focus on this issue.
-For example, FCI (or Fast Causal Inference) algorithm and its variants [61, 62, 63, 64, 65, 66] are proposed to discover causality in the presence of hidden variables.
-And some other related works [67, 68, 69] investigate the causal discovery with soft intervention, that intervenes the SCM but does not change the network structure.
-There are also some related works [70, 71, 72, 73] proposed approaches to search the optimal and efficient adjustment set, that all its variables are observable, minimal and valid, and the removal of any of its variables would destroys the validity.
+For example, the Fast Causal Inference algorithm and its variants [61, 62, 63, 64, 65, 66] are proposed to discover causality in the presence of hidden variables.
+And some other related works [67, 68, 69] investigate causal discovery with soft intervention, which intervenes in the SCM without altering the network structure.
+Related works [70, 71, 72, 73] have also proposed approaches for identifying optimal and efficient adjustment sets, in witch all variables are observable, minimal and valid, and or which removing any variables would invalidate the adjustment.
 4 Differences between POM and SCM
-In this section, the fundamental differences between POM and SCM are discussed, including (i) counterfactual falsifiability: POM assumes the positivity, thus the counterfactuals in POM can be falsified on observations, but SCM does not.
-The counterfactuals in SCM can be outside the observations, thus it is unfalsifiable.
-(ii) using back-door criterion to satisfy ignorability: POM requires ignorability to guarantee the unconfoundedness and correctness of counterfactual inference, but it is a complex human-guided operation.
-By contrast, SCM realizes unconfoundedness by the causal sufficiency and back-door adjustment, that are equivalent to the ignorability assumption.
+In this section, the fundamental differences between POM and SCM are discussed, focusing on three aspects.
+(i) Counterfactual falsifiability: POM assumes positivity; therefore, counterfactuals in POM are, in principle, supported by observations and can be empirically falsified.
+In contrast, counterfactuals in SCM may lie outside the support of the observed data and are thus unfalsifiable.
+(ii) Ignorability and the back-door criterion: POM requires the ignorability assumption to ensure unconfoundedness and the validity of counterfactual inference.
+In practice, satisfying ignorability typically requires identifying appropriate covariates.
+This is a complex, human-guided process.
+By contrast, SCM achieves unconfoundedness through causal sufficiency and back-door adjustment, which are equivalent to the ignorability assumption.
 SCM facilitates counterfactual inference through graphical language.
-(iii) direct causation and indirect causation: POM struggles to process the indirect causation, and it typically relies on Baron-Kenny model.
-But in contrast, SCM has advantages to adjust the front-door path (that has indirect causation) through equivalently applying the back-door criterion twice.
-We further use this advantage to design the forward counterfactual inference algorithm to infer the spatio-temporal counterfactuals autonomously.
+(iii) Direct causation and indirect causation: POM struggles to process the indirect causation, and it typically relies on the Baron-Kenny model.
+In contrast, SCM has the advantage of adjusting for the front-door path (indirect causation) by equivalently applying the back-door criterion twice.
+We further leverage this advantage to design a forward counterfactual inference algorithm that autonomously infers spatio-temporal counterfactuals.
 The following are the details.
@@
 As shown in Fig.
-3, if we query to a POM, “Observed $X_{1}=0.5,X_{2}=1,X_{3}=1.5$, if had done $X_{2}=2$, what would $X_{3}=?$”, we must ensure there are at least two observational samples, and they have the same values $X_{1}=0.5$ but different values in $X_{2}$.
-That is, one is $X_{2}=1$, and another one is $X_{2}=2$, to satisfy the positivity assumption (see Assumption 2.2).
-And then, due to the consistency assumption (see Assumption 2.2), we can infer the counterfactuals through matching the two samples exactly.
-But this is different in SCM, we can infer the counterfactuals through the Pearl’s causal ladder, even if we only have a single sample, as shown in Section 3.1.
+3, consider the following query, “Observed $X_{1}=0.5,X_{2}=1,X_{3}=1.5$, if had done $X_{2}=2$, what would $X_{3}=?$”.
+To answer this query, the positivity assumption requires the existence of at least two observational samples that share the same values $X_{1}=0.5$ but differ in $X_{2}$.
+Specially, one with $X_{2}=1$, and another one with $X_{2}=2$ (see Assumption 2.2).
+And then, under the consistency assumption (see Assumption 2.2), we can infer counterfactuals by exactly matching the two samples.
+In contrast to SCM, we can infer the counterfactuals through Pearl’s causal ladder, even if we only have a single sample, as shown in Section 3.1.
 Thus, how to falsify the counterfactuals?
-It is a difficult question, because the counterfactuals are unobserved according to the definition (see “?” in Table 1), and we need at least two samples for comparison.
-In general, we cannot back to the past to falsify it with a different treatment, and we cannot reproduce the same experiments completely at different times.
+This is a challenging problem, as counterfactuals are, by definition, unobserved (see “?” in Table 1), and we need at least two samples for comparison.
+In practice, we cannot return to the past to apply a different treatment, nor can we perfectly replicate the same experiment at a different time.
 Thus, the inferred counterfactuals in SCM cannot be falsified without the assumptions of positivity and consistency, and this view is rejected ambiguously in Pearl’s book [7].
-Another explanation is that, without the positivity, but with the consistency only in POM, the counterfactuals can be extrapolated by regressive approaches, and they are falsifiable.
+Another explanation is that, without positivity, but with consistency only in POM, counterfactuals can be extrapolated by regressive approaches, and they are falsifiable.
 For example, as shown in Fig.
 5, if there are $T$ and $Y$ only, and the observational data $(T,Y)=\{(0,0.5),(1,1),(2,2.5),(3,2)\}$ is collected, the linear regressive model can be fitted as $Y=0.5T+0.5$.
-Then, we query “if had done $T=4$, what would Y=?Y=?”, the counterfactual answer is $Y=2.5$, and the answer can be falsified by the regressive function because it does not beyond the observation.
-Thus, the consistency assumption actually claims that the counterfactuals can be falsified by the observations.
-This is also true for SCM, because SCM also uses the regressive approaches to infer counterfactuals.
-Moreover, there may be some confusions regarding the preset Bayesian network (see Fig.
-2).
-To build an SCM, a Bayesian network is necessary to be known first.
+We then query “if had done $T=4$, what would Y=?Y=?”.
+The counterfactual answer is $Y=2.5$, and the regressive function can falsify the answer because it does not go beyond the observation.
+Thus, the consistency assumption actually claims that counterfactuals can be falsified by observations.
+This is also true for SCM, which uses regressive approaches to infer counterfactuals.
+Moreover, there may be ambiguity regarding the preset Bayesian network (see Fig.
+3).
+To build an SCM, a Bayesian network must first be known.
 However, how to discover a probabilistic network from a single sample?
 This is usually impossible.
-A reasonable explanation is that, the network structure can be discovered from the system mechanisms (e.g., physical mechanisms, communication connections), or graphical knowledge learned from other domains, instead of probability.
+A reasonable explanation is that the network structure can be inferred from the system’s mechanisms (e.g., physical mechanisms, communication connections) or from graphical knowledge learned from other domains, rather than from probability distribution.
 However, if it is accepted, the definition of causality could be broader and more ambiguous.
-This is why probabilistic and causal Bayesian network are often indistinguishable, as stated in Pearl’s book [7].
-Thus, actually, SCM provides a way to think the counterfactuals, and it emphasizes the thinking way is counterfactual, but not the thinking result, because the result cannot always be falsified without the positivity and consistency assumptions.
-And on the contrary, POM emphasizes the thinking result is counterfactual.
+This is why probabilistic and causal Bayesian networks are often indistinguishable, as stated in Pearl’s book [7].
+Thus, SCM provides a way to think about counterfactuals, emphasizing that the thinking process is counterfactual, not the thinking result, because the result cannot always be falsified without the positivity and consistency assumptions.
+In contrast, POM emphasizes that the thinking result is counterfactual.
 This is a fundamental difference in the problem of counterfactual falsifiability.
+Figure 4: Diagram of data extrapolation.
+Figure 5: Diagram of a back-door path.
 4.2 Ignorability and Back-door Criterion
 In the view of statistics, the ignorability assumption (see Assumption 2.2) is equivalent in calculation to the causal sufficiency assumption (see Assumption 3.1) with a back-door criterion.
-The accurate definition of back-door criterion can be found in these papers and books [17, 18, 4, 5].
-Here we present another from with respect to the Table 1, defined as
+The formal definition of the back-door criterion can be found in these papers and books [17, 18, 4, 5].
+Here we present alternative formulation with respect to Table 1, defined as follows:
 Definition 4.1 (Back-door Criterion [17, 18, 4, 5]).
-In Table 1, the covariate $X$ is said to satisfy back-door criterion, if a Bayesian network is built on all the variables relative to $X,Y,T$, and there is an ordered pair $T\rightarrow Y$ that satisfy
+In Table 1, the covariate $X$ is said to satisfy the back-door criterion, if a Bayesian network is built on all the variables relative to $X,Y,T$, and there is an ordered pair $T\rightarrow Y$ that satisfies
   1. 1.
@@
     $X$ blocks every path between $T$ and $Y$ that contains an arrow into $T$.
-Figure 4: A diagram of data extrapolation.
-Figure 5: A diagram of back-door path.
 Thus, for $X,Y,T$, only one structure can satisfy the back-door criterion, as shown in Fig. 5. And in the back-door path, it is obtained that
@@
 which is the same in case of $do(T=0)$ [17, 18, 4, 5].
-Then, if the causal sufficiency is satisfied, that is, no hidden variable can interfere $X,Y,T$, the expectation of potential outcomes in Table 1 can be calculated by
+Then, if causal sufficiency holds (i.e., no hidden variables confound $X,Y$ and $T$), the expectation of the potential outcomes in Table 1 can be calculated as
 $\mathbb{E}[Y|do(T=1)]$ (9)
@@
 $=$ $\mathbb{E}[\mathbb{E}[Y|T=1,X]].$
-This is the same in POM, as follow:
+This is the same in POM, as follows:
 $\mathbb{E}[Y(1)]$ (10)
@@
 Thus, the ignorability assumption can be decomposed as the causal sufficiency assumption and the back-door criterion, and the causal sufficiency is relatively trivial to satisfy.
-Actually, the ignorability is more likely to be a technical assumption, that needs many controlled trials to deconfound, while the back-door path is clear to search in a Bayesian network.
-A potential and more difficult technical issue is to discover the network structure accurately.
+The ignorability is more likely to be a technical assumption that needs many controlled trials to deconfound, while the back-door path is clear to search in a Bayesian network.
+A potential, more difficult technical issue is accurately discovering the network structure.
 4.3 Direct and Indirect Causation
 Suppose that there is a causal ordered pathway from one treatment variable to another outcome variable.
-Then, if there are one or more variables between the two endpoints, the pathway is called as indirect causation, otherwise, it is called as directed causation.
-Actually, the indirect causation is also called as mediating effect, that has been discussed deeply in Rubin’s papers [74, 75, 76].
-Here, Baron-Kenny model [19, 20] is introduced to present mediating analysis for potential outcome with graph.
+Then, if there is at least one variable between the two endpoints, the pathway is called indirect causation, otherwise, it is called directed causation.
+Indirect causation also called the mediating effect, has been discussed in depth in Rubin’s papers [74, 75, 76].
+Here, the Baron-Kenny model [19, 20] is introduced to present a mediating analysis for a potential outcome using a graph.
 As shown in Fig.
 7, $T\rightarrow Y$ is direct causation, and $T\rightarrow X\rightarrow Y$ is indirect causation.
-Figure 6: A diagram of mediating analysis [19]. Here, $a,b,c$ are coefficients and $U_{X},U_{Y}$ are biases. $b^{\prime}$ and $U_{total}$ are the coefficient and bias respectively for total effect.
-Figure 7: A diagram of front-door adjustment [5]. Here, $T$ is a treatment variable, and $Y$ is a target variable. $X_{1}$ and $X_{2}$ are both covariates.
-To analyse indirect causation, causal-steps approach [19] is proposed, through building three linear regressive models, as shown in Fig.
+To analyse indirect causation, the causal-steps approach [19] is proposed by building three linear regression models, as shown in Fig.
 7.
@@
 Or, to test the significance of $H_{0}:ab=0$ [80, 81].
-If the correct regressive models are built, the counterfactuals can be calculated, just like SCM in Fig.
-2.
-Moreover, they are all regression-based approaches, that restrict the usage in large multivariate datasets, because one needs to test whether a covariate is a confounder or a mediator, and the test process would be time-consuming seriously.
-But on the contrary, the test process can be faster with Bayesian network in an SCM.
-Here, we propose an algorithm to realize forward counterfactual inference as shown in Algorithm 1.
+If the correct regressive models are built, counterfactuals can be calculated, just as in SCM in Fig.
+3.
+Moreover, they are all regression-based approaches that are limited for use in large multivariate datasets, because one needs to test whether a covariate is a confounder or a mediator, and this process would be seriously time-consuming.
+Figure 6: Diagram of mediating analysis [19].
+Here, $a,b,c$ are coefficients and $U_{X},U_{Y}$ are biases.
+$b^{\prime}$ and $U_{total}$ are the coefficient and bias, respectively, for the total effect.
+Figure 7: Diagram of front-door adjustment [5]. Here, $T$ is a treatment variable and $Y$ is a target variable. $X_{1}$ and $X_{2}$ are both covariates.
+However, the test process can be faster with a Bayesian network in an SCM.
+Here, we propose an algorithm for forward counterfactual inference, as shown in Algorithm 1.
 To clearly demonstrate the principle of Algorithm 1, we first introduce the front-door criterion as follows:
 Definition 4.2 (Front-door Criterion [17, 18, 4, 5]).
-The covariate $X$ is said to satisfy front-door criterion, if a Bayesian network is built on all the variables relative to $X,Y,T$, and there is an ordered pair $T\rightarrow Y$ that satisfy
+The covariate $X$ is said to satisfy the front-door criterion, if a Bayesian network is built on all the variables relative to $X,Y,T$, and there is an ordered pair $T\rightarrow Y$ that satisfies
   1. 1.
@@
     all back-door paths from $X$ to $Y$ are blocked by $T$.
-Then, a classic example is provided to show the front-door adjustment through the theorem 3.4.1 in book [5], as also shown in Fig.
+Then, a classic example is provided to show the front-door adjustment using Theorem 3.4.1 in book [5], as shown in Fig.
 7 here.
@@
   1. 1.
-    If there are one or multiple ordered causal pathways from $T$ to $Y$ and there is at least one mediator on the ways, then the causal pathway can be decomposed into two causal pathways through that mediator, e.g., $T\rightarrow X_{1}$ and $X_{1}\rightarrow Y$ in Eq.
+    If there is one or multiple ordered causal pathways from $T$ to $Y$ and there is at least one mediator on the way, then the causal pathway can be decomposed into two causal pathways through that mediator, e.g., $T\rightarrow X_{1}$ and $X_{1}\rightarrow Y$ in Eq.
 (11).
@@
 Thus, the front-door principle is non-atomic and decomposable.
-And with this more foundational principle, we propose the forward counterfactual algorithm to autonomously apply this adjustment principle, as presented in Algorithm 1.
-The algorithm can be conducted on arbitrary variable set $\{X_{1},\dots,X_{n}\}$ that satisfied causal sufficiency.
-And we also provide an example to show the inference process as shown in Fig.
+With this more foundational principle in mind, we propose the forward counterfactual algorithm to autonomously apply it, as presented in Algorithm 1.
+The algorithm can be applied to an arbitrary set of variables $\{X_{1},\dots,X_{n}\}$ that satisfy the causal sufficiency condition.
+We also provide an example to illustrate the inference process, as shown in Fig.
 8.
-In Algorithm 1, the approaches to realize the codes in lines 1 and 3 can be various, e.g., some variants of depth-first-search algorithm and breadth-first-search algorithm [82].
-Thus, their time complexities are both $\mathcal{O}(|\mathbf{V}|+|\mathbf{E}|)$, if a Bayesian network $\mathcal{G}=(\mathbf{V},\mathbf{E})$ is given, where $\mathbf{V}$ is the node set and $\mathbf{E}$ is the edge set.
-Meanwhile, if the computational cost of factorization and adjustment are overlooked, the time complexity of the code in line 9 in Algorithm 1 is $\mathcal{O}(|\mathbf{V}|)$, because we only need to traverse all mediators in each pathway forward, and this process can be performed in parallel.
-Thus, the total time complexity of Algorithm 1 is $\mathcal{O}(|\mathbf{V}|^{2}+|\mathbf{V}||\mathbf{E}|)$.
-Thus, if the network structure of the Bayesian network is large, and it can be discovered accurately, the efficiency of SCM to infer counterfactuals would be better with polynomial time complexity.
-5 Spatio-Temporal Graphical Counterfactuals
-Then for the examples of counterfactual questions at the beginning of this article, that is, if a different investment strategy had been implemented, would we obtain higher returns?
-And, in computer networks, what changes would occur in network load if a node’s configuration had never been changed?
-And, would someone still purchase the product even if they had never been shown the advertisement?
-In these real-world scenarios, interactive behaviors with lagged causal effects are widely-existing.
-POM cannot answer these questions, because POM does not allow the interactive behaviors between experimental units (see SUTVA in Assumption 2.2), even with some temporal quasi-experimental approaches (e.g., Regression Discontinuity Design [83, 84] and Differences in Differences [85, 86]).
-And this is the same to SCM, because the foundational Bayesian network is a DAG that also does not allow mutual connections between network nodes.
-Thus, to answer these counterfactual queries for interactive units and lagged causal effects, a concept of spatio-temporal graphical counterfactuals is proposed here.
 Input: A Bayesian network $\mathcal{G}$ built on $\mathbf{V}=\{X_{1},\dots,X_{n}\}$ with one target $X_{i}$ and multiple sources $X_{j_{1}},\dots,X_{j_{p}}$.
@@
 6    return $P(X_{i}|X_{j_{1}},\dots,X_{j_{p}})$;
+7   end if
 8 else
-8      For every back-door path, select a group of direct predecessors, $\mathbf{W}\subset\mathbf{V}$, and then, according to back-door criterion (see Definition 4.1),
+9      For every back-door path, select a group of direct predecessors, $\mathbf{W}\subset\mathbf{V}$, and then, according to the back-door criterion (see Definition 4.1),
 10    return $\sum_{w}P(X_{i}|X_{j_{1}},\dots,X_{j_{p}},W)P(W)$;
-10
-11
-12Otherwise, for $q$ pathways starting from $X_{j_{p}}$, select $q$ first mediators $X_{j_{p_{1}}},\dots,X_{j_{p_{q}}}$.
+11   end if
+12
+13 end if
+14
+15Otherwise, for $q$ pathways starting from $X_{j_{p}}$, select $q$ first mediators $X_{j_{p_{1}}},\dots,X_{j_{p_{q}}}$.
 Then, a group of source-mediator pairs, e.g., $(X_{j_{p}},X_{j_{p_{q}}})$, can be obtained;
-13Factorize $P(X_{i}|do(X_{j_{1}}),\dots,do(X_{j_{p}}))$ as
+16Factorize $P(X_{i}|do(X_{j_{1}}),\dots,do(X_{j_{p}}))$ as
 $P(X_{i}|do(X_{j_{1}}),\dots,do(X_{j_{p}}))$ (12)
@@
 where $\mathbf{Pa}(X_{j_{p_{q}}})$ is the causal parents of $X_{j_{p_{q}}}$;
-14Solve each distribution $P(\cdot|do(\cdot),\dots,do(\cdot))$ in Eq. (12) by Algorithm 1 recursively.
+17Solve each distribution $P(\cdot|do(\cdot),\dots,do(\cdot))$ in Eq. (12) by Algorithm 1 recursively.
 return $P(X_{i}|do(X_{j_{1}}),\dots,do(X_{j_{p}}))$.
 Algorithm 1 Forward Counterfactual Inference
-Figure 8: An example of Forward Counterfactual Inference algorithm. Here, $X_{i}$ is the same as $X_{i}^{\prime}$ to distinguish two recursions.
+Figure 8: Example of the Forward Counterfactual Inference algorithm. Here, $X_{i}$ is the same as $X_{i}^{\prime}$ to distinguish between the two recursions.
+In Algorithm 1, the approaches to realize the codes in lines 1 and 3 can be various, e.g., some variants of the depth-first-search algorithm and breadth-first-search algorithm [82].
+Thus, their time complexities are both $\mathcal{O}(|\mathbf{V}|+|\mathbf{E}|)$, if a Bayesian network $\mathcal{G}=(\mathbf{V},\mathbf{E})$ is given, where $\mathbf{V}$ is the node set and $\mathbf{E}$ is the edge set.
+Meanwhile, if the computational cost of factorization and adjustment are overlooked, the time complexity of the code in line 9 in Algorithm 1 is $\mathcal{O}(|\mathbf{V}|)$, because we only need to traverse all mediators in each pathway forward, and this process can be performed in parallel.
+Thus, the total time complexity of Algorithm 1 is $\mathcal{O}(|\mathbf{V}|^{2}+|\mathbf{V}||\mathbf{E}|)$.
+Thus, if the network structure of the Bayesian network is large and can be discovered accurately, the efficiency of SCM to infer counterfactuals would be better with polynomial time complexity.
+5 Spatio-Temporal Graphical Counterfactuals
+Then, for the examples of counterfactual questions at the beginning of this article, that is, if a different investment strategy had been implemented, would we obtain higher returns?
+And, in computer networks, what changes would occur in network load if a node’s configuration had never been changed?
+Would an individual still purchase the product had they not been exposed to the advertisement?
+In these real-world scenarios, interactive behaviors with lagged causal effects exist widely.
+POM cannot answer these questions, because it does not allow the interactive behaviors between experimental units (see SUTVA in Assumption 2.2), even with some temporal quasi-experimental approaches (e.g., Regression Discontinuity Design [83, 84] and Differences in Differences [85, 86]).
+This applies to SCM as well, because the foundational Bayesian network is a DAG that also does not allow mutual connections between network nodes.
+Thus, to answer these counterfactual queries for interactive units and lagged causal effects, we propose a concept of spatio-temporal graphical counterfactuals.
+5.1 Spatio-Temporal Bayesian Networks
+Before the SCM, various causal graphical models have been proposed to model the spatio-temporal causality.
+As shown in Fig.
+9, a temporal Bayesian network [59, 87] is proposed to model the momentary causal dependency in a first-order Markov process.
+Further, dynamical Bayesian networks [59] are proposed to represent a combination of multiple temporal Bayesian networks.
+Full time graph [6] is proposed to summarize temporal Bayesian dependencies with complete timestamps.
+Full time graph allows the instantaneous causal effects between two units (or variables) in a high-order Markov process, but it only supports strictly stationary causality, that is, causal dependencies that do not change over time.
+Moreover, the full time graph can be uniquely decomposed into multiple high-order temporal Bayesian networks, named UCN, if the instantaneous causal effects are not allowed [21].
+The comparisons of these causal graphs are shown in Table 2.
+Figure 9: Diagrams of various causal graphs.
 Table 2: Comparisons of various causal graphs.
@@
 Nonstationary causation $\times$ $\checkmark$ $\times$ $\checkmark$
-5.1 Spatio-Temporal Bayesian Networks
-Before the buildings of SCM, actually, various causal graphical models have been proposed to model the spatio-temporal causality.
-As shown in Fig.
-9, Temporal Bayesian Network [59, 87] are proposed to model the momentary causal dependency in a first-order Markov process.
-Further, Dynamical Bayesian Networks [59] are proposed to represent a combination of multiple Temporal Bayesian Networks.
-Full Time Graph [6] is proposed to summarize the temporal Bayesian dependency at full timestamps.
-Note that, Full Time Graph allows the instantaneous causal effects between two units (or variables) in a high-order Markov process, but it only allows strictly stationary causality, that is, the causal dependencies do not change with time.
-Moreover, Full Time Graph is proved to uniquely decomposed into multiple high-order Temporal Bayesian Networks, named UCN, if the instantaneous causal effects is not allowed [21].
-The comparisons of these causal graphs are shown in Table 2.
-Figure 9: Diagrams of various causal graphs.
-Thus, based on the characteristics of these causal graphs, STBNs are defined as a group of DAGs with different timestamps, as shown in Fig.
+Thus, based on the characteristics of these causal graphs, STBNs are defined as a group of DAGs indexed by different timestamps, as shown in Fig.
 9.
-In STBNs, intuitively, the nodes are not equivalent to the variables, but represent the state of the variables at some time steps.
-With the assumptions of causal Markov (see Assumption 5) and faithfulness (see Assumption 6), the joint distribution of all nodes can be factorized as
-$P(X_{1,t},\dots,X_{n,t},X_{1,t-1},\dots,X_{n,t-T+1})$ (13)
-$=$ $\prod_{\tau=0}^{T-1}\prod_{i=1}^{n}P(X_{i,t-\tau}|\mathbf{Pa}(X_{i,t-\tau})),$
-where $X_{1},\dots,X_{n}$ are the variables, and the timestamps are in range of $0\sim T-1$. Theoretically, infinite time steps are allowed in STBNs, but it is usually finite in practice.
-To guarantee the global and unique DAG in STBNs, a temporal assumption is introduced as {assumption}[Temporal Assumption] In STBNs, the cause node precedes or parallels the effect node in time.
+In STBNs, nodes are not equivalent to variables.
+They represent the state of the variables at some time steps.
+Under the assumptions of causal Markov (see Assumption 5) and faithfulness (see Assumption 6), the joint distribution of all nodes can be factorized as
+$P(X_{1,t},\dots,X_{n,t},X_{1,t-1},\dots,X_{n,t-T+1})=\prod_{\tau=0}^{T-1}\prod_{i=1}^{n}P(X_{i,t-\tau}|\mathbf{Pa}(X_{i,t-\tau})),$ (13)
+where $X_{1},\dots,X_{n}$ are the variables, and the timestamps are in the range of $0\sim T-1$.
+Theoretically, STBNs allow an infinite number of time steps, but in practice, they are typically finite.
+To guarantee the global and unique DAG in STBNs, a temporal assumption is introduced as {assumption}[Temporal Assumption] In STBNs, the cause node precedes or occurs concurrently with the effect node.
 This assumption does not allow the directed cause-effect pairs from now to the past, e.g., $X_{i,t-\tau_{1}}\rightarrow X_{j,t-\tau_{2}},\tau_{1}<\tau_{2}$.
-And this guarantees the functionally identifiable network structure without Markov equivalence class [6, 21, 88].
+And this guarantees a functionally identifiable network structure without Markov equivalence classes [6, 21, 88].
 Moreover, the causal sufficiency assumption (see Assumption 3.1) is also needed to guarantee the unconfoundedness.
-Then, if STBNs are given, the spatial-temporal counterfactuals can be inferred through Algorithm 1, because the global and local STBNs are both DAG.
-Compared with the naive Bayesian networks, STBNs have the spatio-temporal concept, and the inferring process is naturally forward along with time.
-This enables to define the momentary interventions for every time steps, and all the interventions would work sequentially, instead of at the same time, like in Algorithm 1.
-If fully consider this property, the inference algorithm can be faster.
+Then, if STBNs are given, the spatial-temporal counterfactuals can be inferred through Algorithm 1, because the global and local STBNs are both DAGs.
+Compared with naive Bayesian networks, STBNs incorporate spatio-temporal concepts, and the inference process naturally proceeds forward in time.
+This enables momentary interventions at every time step, and all interventions work sequentially rather than simultaneously, as in Algorithm 1.
+If this property is considered thoroughly, the inference algorithm can be faster.
 5.2 Nonstationarity of STBNs
-Nonstationarity is a significant topic in researches on time series.
-Actually, however, the reasons inducing it are likely to be different from the view of causality.
-Three of the most currently popular reasons are:
+Nonstationarity is a significant topic in researches on time series. However, the reasons for it are likely to be diff from the view of causality. Three of the most popular reasons are:
   1. 1.
-    Trending: The time series have ascending or descending trending for long terms.
+    Trending: The time series has ascending or descending trending for long terms.
   2. 2.
-    Time-varying Variance: The variances of additive noise are time-varying, that is also called as heterogeneous causality in papers [89, 90].
+    Time-varying Variance: The variances of the additive noise are time-varying, a phenomenon often referred to as heterogeneous causality in the literature [89, 90].
   3. 3.
     Nonstationary Causality: The structures of STBNs change over time, and at least once, as shown in Fig. 10.
-Figure 10: Diagrams of nonstationary causality and heterogeneous causality. Here, $C_{t-\tau}=t-\tau,\tau=0,\dots,T-1$ are usually assumed, thus, they are observable, instead of hidden variable.
-For the trending time series, an effective approach is to calculate differences of time series in preprocessing.
-This approach enables to obtain the differential time series, that is the observations of system dynamics, like Eq.
+Figure 10: Diagrams of nonstationary causality and heterogeneous causality. Here, $C_{t-\tau}=t-\tau,\tau=0,\dots,T-1$ are usually assumed; Thus, they are observable, instead of a hidden variable.
+For trending time series, an effective approach is to calculate the differences of the time series during preprocessing.
+This approach obtains the differential time series, that is, the observations of the system dynamics, as in Eq.
 (14).
-This implied, the dynamics is stable during the period of time, and the structures of STBNs do not change, or in another word, the differential time series are stationary.
+This implies that the dynamics is stable over the period, and that the structures of STBNs do not change, that is, the differential time series are stationary.
 Thus, if the dynamics equations, e.g., Eq.
 (14), are not equal to zero, the causal links in STBNs are identifiable through the Jacobian matrix of dynamics [91, 92, 93, 94].
-For the case of time-varying variance, the causal identifiability is up to the mutually independent exogenous additive noises [88].
-That is, if there is no causal connection between two additive noise nodes, the causation is identifiable.
-Thus, an effective approach is to relate the time-varying noises with a common timestamp, and the timestamps at all time steps are observable confounders $C_{t-\tau},\tau=0,\dots,T-1$ [95, 89, 96, 90].
+In the case of time-varying variance, causal identifiability relies on the mutual independence of exogenous additive noises [88].
+Specially, if no causal connection exists between two additive noise nodes, the causation is identifiable.
+Thus, an effective approach is to relate the time-varying noises with a common timestamp, treating the timestamps at each time step as observable confounders $C_{t-\tau},\tau=0,\dots,T-1$ [95, 89, 96, 90].
 As shown in Fig.
@@
 Then, the noise variables $U_{4,t-1},U_{3,t}$ and $U_{4,t}$ are mutually independent.
-Thus, it is causally identifiable.
-Note that, however, the premise to do so is that any hidden confounders in the case of time-varying variance can be written as smooth or non-smooth functions of timestamps [90].
-Nonstationary causality is challenging but critical, especially for the case of frequently changing structures.
+Thus, causal identifiability is preserved.
+Note that, however, the premise to do so is that any hidden confounders associated with time-varying variance can be expressed as smooth or non-smooth functions of timestamps [90].
+Nonstationary causality presents a significant challenge, particularly in scenarios involving frequently changing structures.
 To identify the nonstationary causality, one approach is to use sliding windows to identify causation within different time periods, as shown in Fig.
 10.
-For each sliding window, it is assumed that the causal structure is invariant, and then, a group of samples would be input to identification algorithms, e.g., [97, 98, 99, 100, 21, 22], to identify a invariant structure during this period.
-However, if the network changes are frequent, the samples used for identification during a short-term period would be few, but in contrast, we usually need sufficient samples to accept the significance of causal connections.
-Thus, another reasonable approach is to identify the change points first, and then, identify the causality with the corresponding periods [101, 102].
+For each sliding window, it is assumed that the causal structure is invariant, and then, a group of samples would be input to identification algorithms, e.g., [97, 98, 99, 100, 21, 22], to identify an invariant structure during this period.
+However, if network changes are frequent, the samples used for identification over a short period would be few.
+In contrast, we usually need sufficient samples to assess the significance of causal connections.
+Thus, another reasonable approach is to first identify the change points, and then, determine the causality for the corresponding periods [101, 102].
 However, this still does not work in the extreme case of frequently changing, because the samples between two change points may be few.
-Moreover, there are some other approaches to view the nonstationary causality as a probabilistic normalized flow, and then use variational Bayesian inference to estimate it globally [103, 104, 105].
+Moreover, there are other approaches that view nonstationary causality as a probabilistic normalized flow, and then use variational Bayesian inference to estimate it globally [103, 104, 105].
 5.3 STBNs and Complex networks
-Here, we suppose STBNs model has been built, and it is associated with interactive time-series process of multivariate (or multiple units).
-Then, we further introduce complex network, another type of network model, to view the causal interactions in complex system.
+Here, we suppose STBNs model has been built, and it is associated with an interactive time-series process of multivariate (or multiple units).
+Then, we introduce a complex network, another network model, to examine the causal interactions in a complex system.
 The network dynamics for variables $X_{i},\dots,X_{n}$ can be defined as
@@
 11.
-Figure 11: A diagram of discretized network dynamics. Here, $A$ is the adjacency matrix of the first-order STBNs, and $A_{ii}^{\prime},i=1,\dots,n$ represents the joined causal effect of $F$ and $G$.
-Thus, the first-order STBNs have some good properties if network dynamics have.
-For example, if we want to know how to affect a node in STBNs by “intervening” another node, the practical inference steps may not be long, due to the small-world property [106, 107] or the scale-free property [108] of complex networks.
-Moreover, the counterfactual outcomes would not change because network systems have tolerance from external “interventions” [109, 110].
+Figure 11: Diagram of discretized network dynamics. Here, AA is the adjacency matrix of the first-order STBNs, and $A_{ii}^{\prime},i=1,\dots,n$ represents the joined causal effect of $F$ and $G$.
+First-order STBNs exhibit advantageous properties when the underlying network dynamics process certain structural features.
+For example, determining how to affect a node in STBNs by “intervening” another node, the practical inference steps may not be long, due to the small-world property [106, 107] or the scale-free property [108] of complex networks.
+Moreover, counterfactual outcomes would not change because network systems have tolerance from external “interventions” [109, 110].
 We would also know the synchronization [111, 112, 113, 114], controllability [115, 116, 117, 118, 119, 120, 121, 122], resilience [23, 123] of STBNs.
-We would also be able to sparsely identify the network structure even if the network size is large [24, 25], because the independence are equivalent to predictability in STBNs [21].
+We would also be able to sparsely identify the network structure even if the network size is large [24, 25], because the independence is equivalent to predictability in STBNs [21].
 Also, the network size of STBNs can be reduced, because many complex networks can be simplified and still provide an insightful description of the causality of interest [124, 125].
-This is meaningful to design counterfactual inference algorithm on smaller networks, and then obtain higher inference efficiency.
+This reduction supports the design of a counterfactual inference algorithm on smaller networks, thereby improving inference efficiency.
 6 Conclusion
-In this work, we mainly focus on the spatio-temporal graphical counterfactuals, and organize an overview for it to discuss its theoretical foundations and application approaches.
-To discuss the theoretical foundations, a survey is investigated, and the definition of counterfactuals is defined based on the concept of potential outcome in the framework of POM firstly.
-And then, the framework of SCM is also introduced to infer counterfactuals through graphical languages, which are equivalent to the counterfactuals in POM.
-Further, to infer counterfactuals on intelligent machine autonomously, a Forward Counterfactual Inference algorithm is designed in this work, and it is able to recursively solve the counterfactual probability distribution, $P(X_{i}|do(X_{j_{1}}),\dots,do(X_{j_{p}}))$, with polynomial time complexity if multi-nodes interventions are conducted on $X_{j_{1}},\dots,X_{j_{p}}$.
-With this algorithm, the spatio-temporal graphical counterfactuals can be inferred if have two elements, that is, the network structure of STBNs and the corresponding algorithm to identify the network structure.
-This work not only discusses the identification algorithms in various nonstationary cases, but also discusses the feasibility to improve the algorithm efficiency from the view of complex networks.
+This work focuses on the spatio-temporal graphical counterfactuals and presents an overview of their theoretical foundations and application approaches.
+To discuss the theoretical basis, a survey is conducted, and the definition of counterfactuals is based on the concept of potential outcomes in the POM framework.
+We then, introduce the SCM framework, which infers counterfactuals via graphical languages equivalent to those in POM.
+Further, to infer counterfactuals on intelligent machine autonomously, a Forward Counterfactual Inference algorithm is designed in this work, and it can recursively solve the counterfactual probability distribution, $P(X_{i}|do(X_{j_{1}}),\dots,do(X_{j_{p}}))$, with polynomial time complexity if multi-nodes interventions are conducted on $X_{j_{1}},\dots,X_{j_{p}}$.
+With this algorithm, spatio-temporal graphical counterfactuals can be inferred, given two elements: the network structure of STBNs and the algorithm identifying it.
+This work discusses structure identification algorithms under various nonstationary conditions and examines the feasibility of improving algorithm efficiency from the perspective of complex networks.
 […]
-This work is supported by the National Key R&D Program of China under Grant No.
-2022ZD0120003, the National Natural Science Foundation of China under Grant Nos.
-62233004, 62273090, 52525204, 52572354, the Jiangsu Provincial Scientific Research Center of Applied Mathematics under Grant No.
-BK20233002, the US National Science Foundation under Grant No.
-2047488 and the Rensselaer-IBM AI Research Collaboration.
+This work was supported in part by the National Key R&D Program of China (Grant No.
+2022ZD0120004), the Youth Scientist Project of the Ministry of Science and Technology of China (Grant No.
+2025YFF0524100), the National Natural Science Foundation of China (Grant Nos.
+62233004, 62273090, 62073076, T2541017, 52525204, 52572354), the Zhishan Youth Scholar Program of the Southeast University, the Jiangsu Provincial Scientific Research Center of Applied Mathematics (Grant No.
+BK20233002), the basic research program of Jiangsu (Grant No.
+BK20253020), the Open Research Project of the State Key Laboratory of Industrial Control Technology, China (Grant No.
+ICT2025B54), the US National Science Foundation under (Grant No.
+2047488) and the Rensselaer-IBM AI Research Collaboration.